### v1

# Why do Machine Learning Notebooks Crash?

Abstract
Jupyter notebooks have become central in data science, integrating code, text and output in a flexible environment. With the rise of machine learning (ML), notebooks are increasingly used for ML prototyping and data analysis. However, due to their dependence on complex ML libraries and the flexible notebook semantics that allow cells to be run in any order, notebooks are susceptible to software bugs that may lead to program crashes. This paper presents a comprehensive empirical study focused on crashes in ML notebooks. We collect 64,031 ML notebooks containing 92,542 crashes from GitHub and Kaggle, and manually analyze a sample of 746 crashes across various aspects, including exception types and root causes. Our analysis highlights unique root causes related to notebook semantics, including out-of-order execution and previous cell error, that have not been thoroughly covered in earlier research. Furthermore, we categorize crashes as ML bugs or general Python bugs and examine how the crashes are distributed across different stages of the ML pipeline. Our analysis reveals that 87% of crashes are caused by API misuse, data confusion, notebook-specific issues, environment problems, and implementation errors. Crashes are more commonly related to ML bugs than general Python bugs, particularly in Kaggle notebooks, where over 70% of crashes are ML-related. The distribution of crashes across ML pipeline stages differs between the two platforms. Additionally, most crashes (58%) occur during data preparation, model training, and evaluation/prediction stages of the ML pipeline. GitHub and Kaggle exhibit different crash distributions across these stages.
Index Terms:
Software bugs, machine learning, Jupyter notebooks, empirical study, crashes
1 Introduction
Computational notebook environments like Jupyter notebooks have become indispensable in data science due to their interactive nature, offering real-time feedback during code execution [1]. This interactivity makes them ideal for tasks such as data visualization, pattern exploration, and data analysis [2, 3]. Notebooks also streamline documentation by combining text, code, execution results, and images into a single, cohesive environment, designed to enhance collaboration and reproducibility [2, 1].
With the growing availability of data, machine learning (ML) and deep learning (DL) techniques have gained popularity for data modeling and decision-making within the broader field of data science [4, 5]. In this context, ML notebooks have been widely adopted for ML prototyping and data analysis [6]. Hence, ensuring the correctness of ML notebooks becomes essential to prevent software bugs from leaking into the later stages of ML development.
ML notebooks are prone to software bugs with various underlying root causes. Such bugs commonly stem from the heavy reliance on ML libraries that are used for various stages in a ML pipeline including data processing, visualization, and building ML models or other data modeling algorithms [7]. These powerful libraries often come with steep learning curves and varying abstraction levels, leading to errors in API usage, configuration, and compatibility. Furthermore, the flexible semantics of Jupyter notebooks, such as allowing code cells to be executed out of order or multiple times, could increase the likelihood of bugs [2, 1, 8]. Understanding these challenges is crucial for improving the debugging process and developing more robust tools for notebook environments. However, most previous studies [9, 7, 10, 11, 12] have been focused on ML scripts rather than ML programs developed in notebooks.
Bugs in ML programs can manifest as low effectiveness, reduced efficiency, or crashes [9] which abruptly terminate the software. In Python, crashes occur as unhandled exceptions, representing the most severe bug symptom as execution is halted and needs immediate attention. Existing research [7, 12, 8] indicates that crashes are the most common symptoms of bugs in ML programs. Crashes are also more easily measurable due to accompanying error information that can be captured in the notebook environment.
We conduct an empirical study to analyze 92,542 crashes in 64,031 ML notebooks by identifying their exception types, root causes, their association with ML libraries, and the ML pipeline stages where crashes typically occur. Further, we highlight root causes related to the improper use of notebook semantics (e.g., out-of-order execution) which has not been covered in previous research. Our goal is to enhance the use of Jupyter notebooks for ML development, guide the creation of better debugging tools, and improve the overall quality of ML notebooks. To our best knowledge, this is the first study focused on analyzing crashes in ML notebooks.
In particular, we examine crashes in notebooks collected from GitHub and Kaggle, addressing three main research questions. Additionally, we investigate how the characteristics of crashes in ML notebooks differ between these two platforms, providing a crosscutting aspect of our analysis.
RQ1: What are the predominant exception types and root causes of crashes in ML notebooks?
Motivation: Crashes in notebooks often provide detailed error information including exception types that describe the direct trigger of the crash. The root cause, on the other hand, is the underlying reason behind a crash such as wrong usage of APIs, mismatched data expectations, or improper use of notebook semantics (e.g., out-of-order execution). Identifying the most common exception types, root causes, and their relationship help developers and researchers prioritize efforts on handling the most frequent issues in a systematic way. This makes the development of ML applications in notebooks more efficient and productive, while developers can still benefit from the flexibility of Jupyter notebooks.
Results: The most common exception types in ML notebooks are NameError (local or global name not found), ValueError (function receives an argument of correct type but inappropriate value), and TypeError (operation or function is applied to an object with wrong type), attributing to 53% of crashes in total. The main root causes include API misuse (19%), notebook-specific issues (19%), data confusion (18%), environment problems (16%), and implementation errors (15%). Notebook-specific crashes, mostly due to out-of-order execution, are a leading cause on GitHub (20%) and third most common on Kaggle (18%). Our analysis of co-occurrences between exception types and root causes reveal that NameError is primarily (56%) caused by notebook-specific factors. API misuse and data confusion are the primary causes of a wide range of exception types.
RQ2: To what extent are crashes in ML notebooks caused by ML bugs compared with general Python bugs?
Motivation: Crashes in ML notebooks can result from ML-related bugs and general Python bugs. We examine to what extent these crashes are caused by ML bugs, including those directly caused by the wrong use of APIs in ML libraries and other errors that arise from the code using ML APIs or variables. This helps guide better ML API usage practices and improves ML library stability.
Results: ML bugs account for 64% of ML notebook crashes, with 59% on GitHub and a significantly higher proportion (71%) on Kaggle. The primary causes of ML-related crashes are API misuse (23%), data confusion (23%), and notebook-specific issues (21%), while general Python bugs are mostly due to environment (24%) and implementation (24%) errors. API misuse is more common in ML-related crashes compared to Python crashes, highlighting the increased difficulty of using ML libraries correctly.
RQ3: In which stages of the ML pipeline are crashes in notebooks most likely to occur?
Motivation: Studying which stages of the ML pipeline are most prone to crashes, such as data preparation, model construction, or model training stages, helps developers and researchers focus on improving robustness in critical stages, and optimize development resources accordingly.
Results: Crashes predominantly occur during data preparation (27%), model training (16%), and evaluation/prediction (15%) stages. GitHub experiences 3.5 times as many crashes unrelated to any ML pipeline stages compared to Kaggle. The two platforms also exhibit distinct crash distributions across pipeline stages. Additionally, over 70% of the crashes during model training and evaluation/prediction are caused by ML bugs. In data preparation, Kaggle has a higher proportion of ML-related crashes compared to GitHub. In contrast, in the environment setup stage, GitHub notebooks show more ML-related crashes.
Replication package. All artefacts can be found at https://github.com/PELAB-LiU/jupyter_nbs_empirical.
Organization. The remaining of this paper is organized as follows. Section 2 provides background information and discusses related work relevant to our study. Section 3 outlines the experimental setup, detailing the methods used for data collection, processing, and analysis. Next, Section 4 presents the results of our study. Section 5 discusses our main findings and addresses their implications and significance. Finally, Section 6 discusses the potential threats to validity of this study and Section 7 concludes the paper.
2 Background and related work
In this section, we describe the concept of a Jupyter notebook and explain the notations used throughout the paper. We then review empirical studies related to bugs in ML and data science programs.
Figure 1: A real Jupyter notebook that trains and evaluates a neural network with documentation (doc cells), Python code (code cells), and execution logs (output and a crash with error information).
TABLE I: Related work of empirical studies on bugs in ML, DL or data science programs, including the considered analysis dimensions, software artifacts, targeted ML libraries, and used data sources: GitHub (GH) and Stack Overflow (SO).
Study Analysis dimensions Targets Libraries Data sources
Zhang et al. [9] Root cause, symptom ML/DL scripts TensorFlow SO posts, GH issues
Islam et al. [7] Bug type, root cause, symptom, ML pipeline ML/DL scripts Caffe, Keras, TensorFlow, Theano, PyTorch SO posts, GH commits
Humbatova et al. [10] Fault taxonomy (root cause) ML/DL scripts TensorFlow, Keras, PyTorch SO posts, GH commits/issues, interviews, survey
Zhang et al. [11] Failure type, root cause ML/DL scripts Not limited Failure messages from a Microsoft DL platform, interviews
Morovati et al. [12] ML/non-ML bugs, root cause and symptom of ML bugs ML/DL systems TensorFlow, Keras, PyTorch GH issues
Ahmed et al. [13] Bug type, root cause, symptom Data analysis programs Most tagged ones on SO SO posts, GH commits/issues
Thung et al. [14], Sun et al. [15], Jia et al. [16], Chen et al. [17] Bug type, root cause, symptom, severity ML frameworks OpenNLP, Paddle, Caffe, TensorFlow, PyTorch, etc GH repositories, bug reports (e.g., JIRA)
De Santana et al. [8] Bug type, root cause, symptom Data science Jupyter notebooks Not limited SO posts, GH commits, interviews, survey
Our study Exception type, root cause, ML/Python bug, ML pipeline ML Jupyter notebooks Not limited Error outputs from GH and Kaggle Jupyter notebooks
2.1 Terminology and motivating example
First we clarify some key terms following the definitions provided in a previous study [18]. We use bugs or faults as inherent defects that result in incorrect or unexpected behaviors, known as errors. If left unaddressed, these errors can lead to failures or crashes. In Python, crashes are represented by exceptions, typically accompanied by error information.
Jupyter notebooks are the most common medium of ML prototyping and data analysis [6]. Notebooks are more flexible compared to, for example, Python scripts by offering additional interactive features. Figure 1 shows an example excerpt of a ML notebook that trains and evaluates a neural network. A notebook typically consists of multiple cells, which can either contain documentation or executable code. When a user runs a code cell, Jupyter uses a kernel to execute it. After execution, the notebook assigns an execution number to the cell, indicating the order in which it was run. The output (if any) is displayed directly below the code cell. In the example, two code cells have been executed, their outputs are shown below, and the execution numbers appear to the left of each cell, marking the execution order.
When an exception occurs during the execution of a code cell (i.e., the code crashes), Jupyter provides detailed error information in the output. The error information includes three key components: exception type, error value, and a traceback that records the sequence of function calls leading to the crash. Analyzing this error information is crucial because it offers direct, first-hand insights into the exact moment and nature of the crash, helping developers understand the immediate cause of failure.
The unique flexibility of Jupyter notebooks introduces additional challenges that may affect how crashes occur. Notebooks allow users to execute cells in any order, run the same cell multiple times, and skip executing cells. While this supports iterative and exploratory development, it can lead to unexpected behavior if cells are executed out of sequence. For example, the notebook in Figure 1 can be executed in reverse, such that the code that evaluates the model is executed before the code that trains the model.
2.2 Studies of bugs in ML or data science programs
Table I summarizes prior empirical research on bugs in ML, DL, and data science programs. It highlights the analysis dimensions, software artifacts, targeted ML libraries, and data sources used in these studies.
The majority of prior research [9, 7, 10, 11, 12] focuses on bugs in ML/DL programs written in scripts. Other studies have examined bugs in popular ML frameworks [14, 15, 16, 17] or data analysis programs [13]. These works have provided valuable insights into the nature of bugs in ML code, helping developers understand common error patterns. The identified root causes include incorrect model parameters or structures, API misuse, and unaligned tensors. Morovati et al. [12] take a different perspective of software maintenance by comparing ML bugs with non-ML bugs, broadening the understanding of bugs in ML/DL systems.
While these studies provide a strong foundation, very few have specifically targeted Jupyter notebooks. One notable exception is the work by De Santana et al. [8], which classifies bugs into eight categories, concentrating on issues related to environment setup, kernel, notebook conversion, and implementation issues. They explore root causes such as configuration and version conflicts, and coding errors, examining their impacts. However, their study does not take into account the iterative and flexible nature of Jupyter notebooks. Furthermore, it primarily examines general bugs without providing an in-depth analysis of those specific to ML. Moreover, no studies focus on program crashes.
Our study addresses these gaps by analyzing crashes in ML notebooks with investigation of how notebook semantics (e.g., out-of-order execution, or previous cell error) impact crashes. Although prior work [2] has linked out-of-order execution to reproducibility issues, we are the first to identify various notebook-specific semantic issues as root causes of crashes. By examining error information in cell outputs, we distinguish ML bugs from general Python bugs and pinpoint the stages of the ML pipeline where these crashes occur.
Most prior research [9, 7, 10, 12, 14, 15, 16, 17, 13, 8] draws from data sources such as Stack Overflow posts, GitHub issues, or code commits. While these sources offer valuable insights, they often describe bugs using natural language, which can abstract away important details about the actual execution context, making it harder to generalize findings. Additionally, Zhang et al. [11] examine DL job failures from an industrial platform. However, job failure messages typically capture issues only after job submission, potentially overlooking critical bugs that occur during coding. In contrast, crash information in Jupyter notebooks provides real-time feedback tied directly to code execution, offering precise and context-rich data. This allows us to capture bugs as they occur, providing more accurate insights into the nature of crashes. By analyzing these crash information, we can identify and diagnose issues more effectively.
Furthermore, many previous studies have focused on bugs in ML/DL programs that use specific ML libraries [9, 7, 10, 13, 12], or defined ML bugs as all errors that occur in ML components in a system [12]. In contrast, our approach is not restricted to a few ML libraries. Instead, we examine a broad range of code from our dataset and identify popular ML libraries. We define ML bugs as bugs in the crashed code that interacts with such libraries, which provides a more direct and precise link between crashes and ML libraries.
Figure 2: Overview of the data collection, filtering, sampling, and manual analysis for Kaggle and GitHub notebooks.
3 Experimental setup
This section outlines the experimental setup of our study, as shown in Figure 2. We begin with the data sources and collection methods, followed by the filtering criteria for identifying the relevant population. Next, we describe the sampling strategy used to select a representative sample. Lastly, we describe the process for manual analysis.
3.1 Data sources and mining method
Our data consists of Jupyter notebooks from GitHub and Kaggle, containing source code, documentation, and output cells (including error information). GitHub is a widely used platform for open-source projects, providing notebooks from a broad community of developers. Kaggle is known for data science and ML competitions, offering notebooks from more specialized ML practitioners. By including both platforms, we capture diverse user bases and practices, allowing us to explore potential differences in their crash behaviors.
GitHub notebooks are downloaded from the deduplicated The Stack dataset [19], which includes active GitHub repositories from January 1, 2015, to March 31, 2022. Kaggle notebooks are retrieved using the KGTorrent database [20], based on the latest (February 7, 2024) Meta Kaggle data [21]. Meta Kaggle is an official and daily updated collection of public metadata of Kaggle notebooks. We download Kaggle Python notebooks with outputs between January 1, 2023, and January 1, 2024. We limit to this time frame because Kaggle limits the number of daily HTTP requests. The time gap between GitHub and Kaggle also indicates fewer overlapping notebooks. This results in a total of 1,070,293 GitHub notebooks and 166,655 Kaggle notebooks.
3.2 Data filtering
We filter the notebooks to identify valid ML notebooks and meaningful crashes, ensuring data relevance and quality.
Python notebooks
This study focuses on notebooks that use Python. To identify the programming language of each notebook, we use the Guesslang library [^1] https://github.com/yoeo/guesslang used in previous works [22, 23] for similar filtering steps. The top three most commonly used languages in GitHub notebooks are Python (95.35%), Julia (1.22%), and R (0.92%). We then filter out non-Python notebooks, leaving a total of 1,020,540 Python GitHub notebooks. All Kaggle notebooks are written in Python.
Notebooks that contain crashes
We aim to study notebook crashes. Hence, we parse and filter all notebooks to retain only those that contain crashes in their cell outputs. Specifically, we filter notebooks containing at least one cell such that 1) its cell_type is code, 2) it comes with output whose output_type is error. After applying this filter, we observe that 10.67% of the collected GitHub notebooks and only 2.61% of Kaggle notebooks contain crashes. This disparity may stem from Kaggle’s role as a platform that encourages developers to share complete, correct, and well-documented notebooks, allowing others to learn from real-world data science challenges and solutions. In total, there are 108,925 GitHub and 4,349 Kaggle notebooks containing 172,638 and 6,745 crashes (i.e., error outputs) respectively.
Valid notebooks
We exclude notebooks with syntactically incorrect code because such errors prevent the code from running at all, making them irrelevant to the actual crashes developers often face. We focus on crashes that occur during code execution, which reveal deeper issues affecting program quality. To check for syntactic validity, we use Python 3 Abstract Syntax Tree (AST) module, which parses code according to Python 3 grammar. Since Python 3 is the current standard, this approach also filters outdated notebooks, aligning with modern ML workflows.
ML notebooks
Jupyter notebooks have various uses. This paper focuses specifically on ML notebooks, defined as those that use at least one ML library. To compile a comprehensive list of ML libraries, we rely on Kaggle as a trusted source, given its reputation as a prominent platform for ML or data science competitions with an active online community. We extract and rank all libraries imported by Kaggle notebooks based on frequency. The top 50 libraries, used by $\sim$95of all Kaggle notebooks, are reviewed by two authors to determine their relevance to ML. From this, we identify 24 popular ML libraries (see Table II), which we use to identify ML notebooks in our dataset.
TABLE II: The 24 most popular ML libraries identified from 166,655 Kaggle notebooks together with their usage frequencies (%), grouped by their main functionality.
Theme ML libraries
Data manipulation & data analysis pandas(82.2%),numpy(81.4%),scipy(8.5%), statsmodels(3.3%)
Data visualization matplotlib(66.0%),seaborn(43.3%),plotly(8.4%), wordcloud(2.1%), missingno(1.6%)
ML frameworks (build/train ML models) sklearn(52.8%),tensorflow(16.4%),keras(7.6%), torch(10.0%),xgboost(8.6%),lightgbm(4.5%), catboost(3.3%),imblearn(2.9%)
Computer vision cv2(7.7%),torchvision(4.3%),skimage(1.5%)
NLP nltk(4.3%),transformers(3.7%)
Optimization & model tuning optuna(1.5%)
Dataset handling datasets(1.4%)
KeyboardInterrupt exceptions
A KeyboardInterrupt exception happens when a user manually stops program execution (e.g., with Ctrl+C). We exclude notebooks containing only KeyboardInterrupt type of exceptions because it is user-triggered and does not necessarily reflect code issues.
Enriching and filtering non-descriptive error values
Normally, error values of crashes can be extracted from notebooks without extra processing. But, in some cases, developers substitute error values with non-descriptive messages like ignored or null. We replace them with a more descriptive message extracted from the traceback of the crash. When this is not possible, the respective crash is removed from the dataset.
The final population comprises 61,342 GitHub notebooks with 88,667 crashes and 2,689 Kaggle notebooks with 3,875 crashes, in total 64,031 notebooks with 92,542 crashes.
3.3 Sampling
Our study requires a manual labeling and review process to analyze ML notebook crashes. Since investigating the entire population is infeasible, we select a meaningful sample of crashes for manual analysis similarly to De Santana et al. [8].
We apply proportional stratified sampling [24], a strategy that builds on top of simple random sampling. This method divides the population into strata and randomly samples from each group in proportion to its size. We base these strata on error values as they capture more detailed information of crashes compared to, for example, exception types alone. Additionally, crashes with different exception types but similar error values often exhibit similarities, suggesting they share common root causes or other patterns.
Clustering crashes
To group similar error values of crashes, we apply a clustering method. We first preprocess error values to generalize them, such as removing unnecessary substrings like URLs, file paths, function names, variable names, and punctuation. This makes the error values more representative of common crash patterns rather than overly specific to individual crashes.
We then cluster the preprocessed error values using Jaccard similarity [25] which measures overlap between sets. This method is ideal for grouping error values with similar patterns, as shown in previous studies [26, 27]. An example of clustering results using Jaccard similarity is presented in Table III. We chose a threshold of 0.7 to perform Jaccard similarity based on preliminary experiments. We first cluster all crashes (both GitHub and Kaggle) to get a comprehensive view of common patterns, then calculate clusters for GitHub and Kaggle datasets separately to identify platform-specific differences in crash behaviors. Each pair of error values is compared, and if their similarity exceeds the threshold, they are connected. Crashes are grouped into the same cluster if they can be reached through any connected crashes, ensuring reachability within the cluster. We obtain a total of 7,433 clusters. Separate statistics for each platform reveal 7,222 clusters for GitHub and 640 for Kaggle.
TABLE III: Jaccard similarity result of two error values.
Error value (original) Error value (processed) Jaccard similarity
name ’x’ is not defined name is not defined 1.0 (the same)
name ’head’ is not defined name is not defined
We further visualize the clusters in relation to exception types. We observe that multiple exception types have similar error values as Table IV shows. This is likely because Python exceptions can be customized and raised by users or library developers under various names or reasons. As a result, exceptions with different names may represent similar issues. This finding motivates our approach of clustering based on error values, rather than using exception types directly.
We include all clusters, emphasizing the most frequent ones, which cover over 70% of crashes in each dataset - 65 clusters for GitHub with a minimum size of 100 and 50 for Kaggle with a minimum size of 10. The remaining, less common crashes are grouped into a single cluster. This results in 66 clusters for GitHub and 51 for Kaggle overall.
TABLE IV: Top 5 clusters across GitHub and Kaggle datasets. The first column shows cluster sizes (CS) with the number of exception types (#ET) per cluster. The second column lists the top 3 exception types and an example error value. Python built-in exception types are explained, while customized exceptions are marked as “self-defined.”
CS (#ET) Top 3 ET (exception types) Example error value
25070 (3) NameError(25050): local/global name not found UndefinedVariableError(19): self-defined RemoteError(1): self-defined
name ’sc’ is not defined
6107 (6) AttributeError(6072): attribute reference/assignment fails ModuleAttributeError(18): self-defined TypeError(13): operation/function apply to wrong type object
’numpy.ndarray’ object has no attribute ’sample’
5136 (3) ModuleNotFoundError(4568): module could not be located ImportError(567): cannot load a module DependencyNotinstalled(1): self-defined
No module named ’src.utils.compute_overlap’
4620 (83) KeyError(3911): a mapping key is not found SystemExit(349): raised by the sys.exit() ValueError(48): argument with right type but wrong value
’Hero Over’
3410 (16) FileNotFoundError(3230): file/directory not exist IOError(79): aliases of OSError OSError(28): system-related error,  e.g., ’disk full’
No such file or directory: ’GloVe/glove.840B.300d.txt’
TABLE V: Classification dimensions of crashes in ML notebooks.
RQ Dimension Category Description
RQ1 exception type Python built-in exception types (e.g., NameError, ValueError, TypeError, KeyError) These are exceptions that occur during the execution of Python code, typically caused by issues like undefined variables, incorrect data types, or illegal key errors. Some examples can be seen in Table IV, more details can be found in the Python documentation.
others (e.g., APIArgumentError, EnvironmentError, ResourceError) These consist of library customized exceptions. For example, APIArgumentError arise when argument values fail to meet expected constraints of an API, EnvironmentError include import exceptions and system configuration issues, ResourceError contain exceptions such as out of memory or out of disk space.
root cause data confusion (DATA) Errors occur when the expected data structure (e.g., shapes, value types, feature names) does not match the actual input, or previously saved ML models (considered as external data) fails to align with the model defined in the code (e.g., model structure).
API misuse (API) Using an API incorrectly, e.g., providing wrong arguments or applying an API improperly.
notebook-specific issues (NB) Errors that stem from the unique execution behavior of notebooks, e.g., out-of-order execution, and unresolved errors in prior cells (previous cell error).
implementation error (IMPL) Faults in the coding logic, e.g., incorrect implementation or missing appropriate conditions.
library change (LIB) Errors caused by updates to an API or libraries, making current code incompatible.
environment issues (ENV) Issues due to incorrect environment setting, e.g., missing/incompatible libraries, file path errors.
intentional (INTENT) Errors deliberately introduced by users, often for testing or education purposes.
insufficient resources (RSC) Crashes due to resource constraints, like running out of memory or disk space.
unknown (UNK) Errors for which the root cause cannot be determined.
RQ2 ML/Python bug ML bug Errors in code using ML libraries or related components.
Python bug Errors from general Python code, unrelated to ML libraries.
RQ3 ML pipeline environment setup (ENVS) Errors occurring when setting up the environment, e.g., missing modules or configuration issues.
data preparation (DATAP) Errors occurring during data preprocessing or preparation for analysis and training.
data visualization (DATAV) Errors encountered while visualizing data.
model construction (MCONS) Issues related to building, compiling, or visualizing ML models.
model training (MTRAIN) Errors that arise during model training, including hyperparameter tuning and validation.
evaluation/ prediction (EVAL) Errors related to evaluating model performance or making predictions.
no ML pipeline (NONE) Errors unrelated to ML pipeline (e.g., notebooks for tutorials, education, or simulations).
Proportional-to-cluster-size sampling
We estimate sample sizes using random sampling without replacement [28, 29]. We calculate sample sizes and analyze sampled crashes separately for GitHub and Kaggle because they may exhibit different behaviors. The desired sample sizes are determined with a 95% confidence interval (CI) and a 5% margin of error (MoE), aiming for results that are 95% accurate within a 5% margin of error from the entire population. This results in sample sizes of 360 for GitHub and 332 for Kaggle crashes.
We then use these sample sizes and the identified clusters (GitHub: 66, Kaggle: 51) to conduct proportional sampling based on cluster size. To ensure comprehensive coverage, we sample at least one crash from each cluster, as recommended by Chan et al. [24]. This yields a total of 390 sampled crashes from GitHub and 356 from Kaggle.
3.4 Crash analysis
We manually analyze crashes from various perspectives. We establish a set of dimensions to describe crashes in ML notebooks. Table V shows the 4 dimensions considered in our analysis and how they relate to our research questions.
EXCEPTION TYPE
The first dimension categorizes crashes based on the exceptions that trigger them, originally shown as the error names in the notebook outputs (see Figure 1). This classification is directly linked to the specific runtime exceptions responsible for a crash. A comparable approach is used by Zhang et al. [11], who classified job failures, which are instances where the job crashes, according to their runtime error types.
We manually refine the raw exception types for two reasons. First, we distinguish between Python’s built-in exceptions and those defined by third-party libraries. Our analysis reveals that 7.21% of crashes on GitHub and 7.31% on Kaggle have names differing from Python’s built-in types, indicating these exceptions are likely created by third-party libraries or users. By refining exception types, we can better identify library-specific exceptions. Moreover, as shown in Table IV, different exception types may exhibit similar patterns, while the same exception type may represent different triggers. Users can define and raise exception types freely, which may not always accurately represent the true trigger of a crash. Therefore, manually refining these types allows us to enhance our analysis of the crashes.
ROOT CAUSE
This dimension categorizes root causes of crashes, focusing on the fundamental issues that lead to these crashes. Unlike traditional software, ML programs face unique challenges, such as complex data processing, model construction, and extensive use of ML APIs. The flexibility of notebooks, which allows cells to be executed at any time, can introduce further complications.
We base our initial root cause categories on prior studies [7, 11, 8]. We also explore how notebook semantics, such as out-of-order execution and unresolved errors in previous cells, contribute to crashes. By categorizing crashes by root causes, we capture a broad range of crash sources, from misunderstanding of input data or APIs to environment-specific and notebook-specific issues, obtaining a comprehensive understanding of why the crashes happen.
ML/PYTHON BUG
This dimension classifies crashes based on whether they occur in code related to ML libraries. Morovati et al. [12] compare ML bugs and non-ML bugs by identifying ML bugs as all bugs within ML components of DL systems. In contrast, our study focuses on the ML components themselves, aiming to evaluate how often crashes stem from direct interactions with ML libraries. We define ML bugs as crashes that occur in code using ML libraries or objects derived from them. By distinguishing ML bugs and general Python bugs, we can better understand the challenges unique to ML development, including the use of ML APIs, model handling, and data processing tasks.
ML PIPELINE
In this dimension, we categorize crashes based on where they occur during the ML pipeline to identify critical stages. Building on the framework from Islam et al. [7] and a pilot study of our dataset, we define key stages including environment setup, data preparation, data visualization, model construction, model training, and evaluation/prediction. This dimension is crucial for analyzing where in the ML pipeline crashes are more frequent and potentially harder to resolve.
3.5 Manual labeling process
To classify ML notebook crashes, our process is structured in two stages: labeling and reviewing.
In the labeling stage, we categorize crashes across the predefined dimensions introduced in Section 3.4. We adopt grounded theory methods [30], incorporating first-cycle and second-cycle coding phases [31]. We begin with an initial set of categories for each dimension, derived from prior studies [11, 7, 12] and our pilot study, and iteratively refines them as needed to maintain classification accuracy.
During the first-cycle coding, three evaluators independently label an initial set of 50 crashes from the sampled dataset. Following this, the evaluators meet to synchronize their findings, discuss the results, and establish uniform labeling principles. This collaborative discussion promotes consistency and mitigates potential bias early in the process.
To further improve consistency, we organize crashes from both GitHub and Kaggle datasets by clusters, which groups similar crashes together for labeling and review. Each crash is categorized based on the predefined dimensions, with new categories introduced only when necessary. Additionally, evaluators reference all notebooks containing the sampled crashes to access relevant crash contexts. This structured approach, grounded in an open coding scheme [30], allows category expansion while minimizing bias and enhancing consistency in our classifications.
During the second-cycle coding, the remaining crashes are labeled independently. Regular meetings are held to discuss and reach consensus on any disagreements, promoting consistent standards across the entire dataset.
In the reviewing stage, a separate evaluator conducts reviewing to validate the labeling results. In cases of disagreement, discussions take place to resolve conflicts, and if needed, all three evaluators come together to reach a final decision. The change rates before and after reviewing are shown in Table VI. The process sometimes results in updates to earlier categories, particularly as the classification standards are refined during review. For example, many changes are made to categorize crashes in ML pipeline based on the stage where the code is written, rather than by its functionality, which leads to a high rate of reclassification. On the other hand, the high change rate of root cause reflects the challenges of accurately identifying the root causes of the crashes.
TABLE VI: Changes of labeled crashes after reviewing.
RQ Dimension Changed total Percentage
RQ1 exception type 37 14.51%
root cause 81 31.76%
RQ2 ML/Python bug 25 9.80%
RQ3 ML pipeline 82 32.16%
After manually labeling and reviewing, we find four irrelevant crashes, such as cells containing random text like “end,” or exceptions raised to print comments like “Done!” These do not represent actual crashes and are excluded from our analysis, resulting in 389 labeled crashes from GitHub and 353 from Kaggle. Despite this reduction, the sample sizes remain statistically significant with 95% confident level.
4 Results
Next, we present the results to answer each research question. To investigate correlations between two distributions, we use the $\chi^{2}$ test to assess whether there is a significant difference, with a null hypothesis stating no difference. We adopt a 95% confidence level, meaning that a $p$-value less than 0.05 indicates a significant difference. For distributions with small sample sizes ($<5$), we apply Barnard’s exact test [32], using Bonferroni correction [33] when the categories are non-binary.
4.1 RQ1: What are the predominant exception types and root causes of crashes in ML notebooks?
(a) Exception types
(b) Root causes (see Table V for the full names)
Figure 3: Overview of the exception types and root causes of crashes in Jupyter notebooks from GitHub and Kaggle.
Crashes in ML notebooks stem from various types of exceptions. We refine and enhance the raw exception types typically shown in the crash information. Additionally, we complement these classes by identifying their underlying root causes and discussing how they relate to exception types. The results are visualized in Figure 3 and Figure 4.
Exception types
3(a) highlights the top 10 most frequent exception types. The most common exception type is NameError, accounting for 30% of cases (GitHub: 31%, Kaggle: 28%), followed by ValueError at 12% (GitHub: 11%, Kaggle: 12%), TypeError at 11% (GitHub: 10%, Kaggle: 12%), IOError at 9% (GitHub: 8%, Kaggle: 11%), and APIArgumentError also at 9% (GitHub: 10%, Kaggle: 8%).
The $\chi^{2}$ test shows no significant difference in exception type distributions between GitHub and Kaggle ($p>0.05$).
Finding 1.1: The most common exception types of ML notebook crashes are NameError (30%), ValueError (12%), and TypeError (11%).
Root causes
3(b) visualizes the distribution of root causes. The most common ones are API misuse (19% overall; GitHub: 18%, Kaggle: 21%), notebook-specific issues (19% overall; GitHub: 20%, Kaggle: 18%), data confusion (18% overall; GitHub: 17%, Kaggle: 19%), environment issues (16% overall; GitHub: 17%, Kaggle: 16%), and implementation errors (15% overall; GitHub: 15%, Kaggle: 15%).
Finding 1.2: The main root causes of ML notebook crashes include API misuse (19%), notebook-specific issues (19%), data confusion (18%), environment problems (16%), and implementation errors (15%).
Notebook-specific issues are a major cause of crashes, ranking first on GitHub and third on Kaggle. The majority of these crashes result from out-of-order execution (GitHub: 64%, Kaggle: 75%), where cells are executed in an incorrect sequence. This occurs when users run notebook cells without considering dependencies. Another common cause is previous cell error (GitHub: 36%, Kaggle: 25%), where users encounter crashes in earlier cells but continue executing dependent cells without resolving the issues, resulting in additional crashes.
The root cause distribution does not differ significantly between GitHub and Kaggle ($p>0.05$).
Finding 1.3: Notebook-specific issues are the leading cause of crashes on GitHub ML notebooks and the third most common on Kaggle, with out-of-order execution being the most frequent (GitHub: 64%, Kaggle: 75%) problem.
Root causes of exception types
We merge both sample sets (GitHub and Kaggle) and plot the co-occurrence of exception types and their root causes in 4(a).
Interestingly, we observe that root causes like API misuse, data confusion, and implementation errors contribute to a wide range of exceptions in ML notebooks. Notably, API misuse alone is responsible for at least 20% of crashes across multiple exception types, including ValueError (28%), TypeError (32%), APIArgumentError (63%), AttributeError (41%), KeyError (20%), and IndexError (27%). Meanwhile, data confusion contributes to over 30% of ValueError (43%), TypeError (35%), KeyError (55%), and IndexError (54%). These findings indicate that developers need to be attentive to the proper use of APIs and ensure that their data inputs are correctly formatted.
Finding 1.4: API misuse and data confusion are the primary causes of a broad range of exception types.
(a) Co-occurrence of exception types and root causes.
(b) The percentage point (pp) difference, such that GitHub has $x$ pp more/fewer instances than Kaggle.
Figure 4: Overview of the co-occurrence of crash exception types and root causes (see Table V for full names), and the difference in co-occurrence between Kaggle and GitHub.
Certain exception types are dominantly linked to a specific root cause. For instance, IOError is connected to environment issues, APIArgumentError to API misuse, and KeyError and IndexError to data confusion. However, exceptions like ValueError, TypeError, and AttributeError have a wider range of root causes, indicating that these crashes may arise from more complex interactions in the code. Specifically, ValueError and TypeError can arise from data confusion (43%, 35%), API misuse (28%, 32%), and implementation errors (12%, 16%). AttributeError is primarily caused by API misuse (41%), implementation errors (20%), and data confusion (16%).
Moreover, we find that 56% of NameError exceptions stem from notebook-specific issues which also account for 20% of KeyError occurrences. This highlights how the unique notebook semantics can introduce distinct challenges that manifest in particular exception types.
Finding 1.5: ValueError, TypeError, and AttributeError arise from diverse root causes. NameError is mainly due to notebook-specific issues.
4(b) highlights the absolute percentage differences in co-occurrence between GitHub and Kaggle notebook crashes. Barnard’s exact tests with Bonferroni correction show no significant difference in root cause distributions between GitHub and Kaggle for all exception types.
However, some differences appear in the sample. Data confusion and API misuse both cause TypeError 1.6 times more commonly on Kaggle than on GitHub (a percentage points (pp) difference of 16). On GitHub, these are 4.5 and 1.8 times more commonly caused by intentional bugs and implementation errors (16 and 9 pp difference, respectively). Moreover, KeyError crashes are 1.5 times more commonly caused by data confusion on GitHub (23 pp difference), whereas notebook-specific issues are 3 times more frequent on Kaggle (18 pp difference).
4.2 RQ2: To what extent are crashes in ML notebooks caused by ML bugs compared with general Python bugs?
Crashes in GitHub and Kaggle notebooks can stem from either code that uses ML libraries or general Python code. We analyze the distribution of ML/Python bugs to understand ML-specific crashes. Additionally, we examine the root causes of both ML and Python bugs, as shown in Figure 5, to gain deeper insights into their distinctions and similarities.
ML bugs and Python bugs
ML bugs account for 64% of crashes overall in ML notebooks, with 59% on GitHub and 71% on Kaggle. The $\chi^{2}$ test reveals a significant difference in the distribution of ML and Python bugs between the two platforms ($p<0.001$). These findings suggest that a substantial proportion of crashes on both platforms stems from code that directly interacts with ML libraries, with the impact being particularly pronounced in Kaggle notebooks.
Finding 2.1: A higher percentage of crashes in both GitHub and Kaggle notebooks are attributed to ML bugs (GitHub: 59%, Kaggle: 71%) compared to Python bugs. Kaggle exhibits an even more significant prevalence of ML-related crashes than GitHub.
Root causes of ML and Python bugs
As Figure 5 illustrates, for ML-related crashes, the primary causes are API misuse (23%), data confusion (23%), and notebook-specific issues (21%). Environment setup problems (12%) and implementation errors (10%) are also big factors. In contrast, general Python crashes are mainly driven by environment issues (24%) and implementation errors (24%), followed by notebook-specific issues (17%) and API misuse (13%).
Finding 2.2: The primary causes of ML-related crashes are API misuse (23%), data confusion (23%), and notebook-specific issues (21%), while general Python crashes are mostly caused by environment problems (24%), implementation errors (24%), and notebook-specific issues (17%).
Figure 5: Overview of the overall co-occurrence of ML/Python bugs and root causes (see Table V for full names).
The root cause distributions between ML and Python bugs differ significantly ($p<0.001$). Notably, the higher percentage of API misuse in ML-related crashes (23%) compared to general Python crashes (13%) underscores the challenges developers face in correctly using ML libraries.
We find no statistically significant differences between the platforms in the distribution of root causes for either ML bugs or Python bugs ($p>0.05$ for both).
Finding 2.3: API misuse is more prevalent in ML-related crashes (23%) than in general Python crashes (13%).
4.3 RQ3: In which stages of the ML pipeline are crashes in notebooks most likely to occur?
The ML pipeline outlines key stages of ML programs. We analyze the frequency of crashes occurring in these stages during ML development in notebooks and determine how often crashes in a certain stage involve ML bugs versus general Python bugs. The results are shown in Figure 6.
ML pipeline
6(a) shows that most crashes occur during the data preparation (27% overall; GitHub: 29%, Kaggle: 26%) stage, model training (16% overall; GitHub: 13%, Kaggle: 19%), evaluation/prediction (15% overall; GitHub: 11%, Kaggle: 21%), and data visualization (11% overall; GitHub: 10%, Kaggle: 12%) stages.
Finding 3.1: The majority of crashes occurs during data preparation (27%), model training (16%), and evaluation/prediction (15%) stages in the ML pipeline.
Notably, 25% of crashes on GitHub are unrelated to any pipeline stage, which is approximately 3.5 times more than Kaggle (7%). We apply the $\chi^{2}$ test to evaluate this, revealing a significant difference in distributions ($p<0.001$).
(a) ML pipeline stages
(b) Adjusted residuals
Figure 6: Overview of crashes across the ML pipeline (see Table V for full names) with the distribution of ML/Python bugs per stage and the adjusted residuals per stage.
Finding 3.2: GitHub notebooks show 3.5 times more crashes than Kaggle not related to ML pipeline stages.
We then examine differences between GitHub and Kaggle in ML pipeline stages, excluding crashes unrelated to any stage. A $\chi^{2}$ test reveals a significant difference ($p<0.01$). To understand which stages contribute more to this difference, we calculate the adjusted residuals [34], comparing observed crashes in each stage with the expected number of crashes under the null hypothesis (shown in 6(b)). A larger absolute residual indicates a greater contribution. The results suggest that GitHub contains more crashes in the environment setup and data preparation, and Kaggle in the evaluation/prediction stage.
Finding 3.3: GitHub and Kaggle crashes are distributed differently across ML pipeline stages, especially in environment setup, data preparation, and evaluation/prediction.
ML/Python bugs per ML pipeline stage
The breakdown of ML and Python bugs in each pipeline stage reveals that ML bugs dominate critical stages, with over 80% of crashes during model training and construction and over 70% during evaluation/prediction for both platforms.
Finding 3.4: ML-related crashes are dominant (over 70%) in model training, construction, and evaluation/prediction stages of the ML pipeline.
We apply Barnard’s exact test to assess significant differences of ML/Python bug distributions between GitHub and Kaggle for each ML pipeline stages. The results show significant differences in the environment setup ($p<0.05$) and data preparation ($p<0.01$) stages. In Kaggle, 70% of crashes during data preparation are ML-related, whereas in GitHub, ML and Python bugs occur at similar rates. In the environment setup stage, most Kaggle crashes are due to Python bugs, only 16% are ML bugs, while in GitHub, 45% of crashes involve ML bugs.
Finding 3.5: Kaggle has 1.4 times more ML-related crashes in the data preparation stage than GitHub, whereas GitHub shows almost 3 times more in the environment setup stage.
Root causes of ML pipeline stages
We further analyze root cause distribution across ML pipeline stages (see Figure 7). Most stages display a range of root causes, including API misuse, notebook-specific issues, data confusion, environment issues, and implementation errors. Notably, in the environment setup stage, 84% of crashes stem specifically from environment-related issues.
Figure 7: Overview of the overall co-occurrence of ML pipeline stages and root causes (see Table V for full names).
5 Discussion
This section discusses the key implications of our findings and the differences between GitHub and Kaggle platforms. We then compare our findings with existing studies.
5.1 Practical implications
Notebook-specific issues
As discussed in RQ1 (Section 4.1), notebook-specific issues are a major cause (19%) of crashes, contributing to 56% of NameError and 20% of KeyError exceptions. This highlights how notebook semantics can become a liability, increasing the risk of crashes. For instance, if a variable x is defined in one cell but a later cell that references x is executed first, a NameError occurs due to out-of-order execution since x has not been initialized yet. Similarly, if a developer incorrectly tries to add a column to a dataframe, causing the cell to crash, and then runs another cell that references this non-existent column without first correcting the initial error, a KeyError will occur due to previous cell error. The flexibility to execute cells in any order can easily disrupt cell dependencies, especially in complex ML notebooks. This underscores the unique challenges of bug detection and debugging in notebooks, which requires careful tracking of variable states and cell execution order at the current execution point. NBSAFETY [35] addresses this issue by providing a custom Jupyter kernel that flags potentially unsafe cell executions.
Takeaway 1: The flexibility of the notebook environment can cause instability, as out-of-order execution often leads to crashes like NameError.
Data confusion
Another common root cause of crashes is data confusion (18%), indicating practical challenges in handling heterogeneous data types. It is a key contributor to a variety of exceptions such as ValueError, TypeError, KeyError, and IndexError, indicating that tackling data confusion could help resolve many of these exceptions. Data confusion also accounts for a high percentage of the ML-related crashes (23%), as discussed in RQ2 (Section 4.2). It suggests that improper data handling such as incorrect data manipulations for ML APIs, is critical in ML-specific code. Specifically, ML development often requires transformations of data into specific structures that can be easily mishandled, causing errors like shape mismatches, improper data transformations, or incorrect assumptions about data formats. This underscores the need for improved data validation.
Takeaway 2: Data handling poses a big challenge when using ML notebooks.
ML bugs
When analyzing the results for RQ2 (Section 4.2), we find that crashes in ML notebooks are overall 1.8 times more frequently linked to ML bugs compared to Python bugs, reflecting key differences between developing ML and traditional software. Notably, a significantly higher proportion of ML-related crashes (23%) are directly attributed to API misuse, compared to only 13% for general Python crashes. The extensive reliance on ML libraries increases API misuse errors, particularly when developers are unfamiliar with libraries’ specific requirements. Although these libraries are crucial for effective data processing and modeling, they often have complex usage guidelines [36, 37]. Misinterpretations of parameter settings, or input-output expectations frequently result in crashes. To mitigate these issues, methods that verify API usage and enhance error-handling mechanisms could be beneficial.
Takeaway 3: Correct use of ML libraries is challenging in ML notebook development, highlighting the importance of verifying API usages and enhancing error-handling.
The data preparation stage
RQ3 (Section 4.3) demonstrates that most ML notebook crashes occur during the data preparation stage (GitHub: 29%, Kaggle: 26%) of the ML pipeline. This highlights the primary challenge developers face: preparing or preprocessing datasets to ensure compatibility with model requirements. Issues often arise from understanding the data itself or aligning it with API expectations. Additionally, the limited support available for developers during this stage implies they may spend a disproportionate amount of time here. These insights suggest that improving tooling and resources for data preparation could significantly enhance developer productivity.
Takeaway 4: Data preparation is the most crash-prone stage in ML pipelines, highlighting the necessity for enhanced support to help developers navigate this critical phase.
The model training and evaluation/prediction stages
Crashes are more common in the model training (16%) and evaluation/prediction (15%) stages than in the model construction (7%) stage. These errors typically arise at runtime, when models are trained or evaluated with actual data inputs, rather than during the construction phase when input shapes are often flexible. Libraries like TensorFlow, for instance, may not fully validate data compatibility until training begins, which can lead to errors surfacing only in these later stages. When issues occur within specific model layers, error messages can be ambiguous, making it difficult for developers to pinpoint and resolve problems. Consequently, developers may need to iteratively debug by re-training and re-evaluating multiple times, a tedious and time-intensive process.
Notebook semantics help preserve previous cell executions, so developers can avoid re-running all preceding cells (e.g., data preparation). However, this approach also increases the risk of crashes from, for example, out-of-order execution, where dependencies may no longer align. These challenges highlight the importance of enabling early bug detection without requiring code execution for training or evaluating ML models, along with mechanisms to track cell dependencies and maintain the correct execution order.
Takeaway 5: Early bug detection through monitoring cell execution order in notebooks would catch issues before code execution, minimizing repetitive debugging cycles.
5.2 Comparison between GitHub and Kaggle
Crash rate
When filtering notebooks with crashes, we find that 10.67% of GitHub Python notebooks contain error outputs, compared to 2.61% on Kaggle. This difference likely reflects GitHub’s broader use, such as version control and backup, where users upload more experimental and incomplete code, while Kaggle encourages sharing more complete solutions for ML competitions and tasks.
ML pipeline
Comparing ML notebook crashes between GitHub and Kaggle reveals their key differences. GitHub has a significantly higher proportion of crashes outside typical ML pipeline stage (25% vs. 7% on Kaggle), suggesting a more diverse use of notebooks where crashes often occur outside standard ML workflows. Moreover, there is a notable difference in crash distributions across ML pipeline stages between the two platforms. For example, GitHub has more crashes during environment setup, likely due to its user-driven environments that can lead to inconsistencies in software versions or dependencies. GitHub also sees more crashes in data preparation, likely because it often uses “raw” datasets that require more complex preprocessing, whereas Kaggle provides standardized datasets with detailed descriptions. In contrast, Kaggle has more crashes during evaluation/prediction, possibly due to its submission process for competitions, which introduces additional model evaluation complexities.
ML/Python bug
Kaggle shows a significantly higher prevalence of ML-related crashes than GitHub (71% vs. 59%). One exception is during the environment setup stage of the ML pipeline, where GitHub shows 3 times more ML-related crashes. This is likely because Kaggle’s pre-configured environment includes essential ML libraries, reducing setup-related errors. In contrast, GitHub, as a general-purpose code repository, preserves crashes from various environments. During the data preparation stage, Kaggle notebooks face 1.4 times more ML-related crashes, suggesting that Kaggle users focus more on ML-specific tasks, while GitHub users encounter a broader range of coding challenges. Furthermore, Kaggle provides a wealth of datasets alongside its notebooks, making it a valuable resource for reproducing experiments. However, existing studies [2, 38, 1, 6, 8] on notebooks have largely overlooked it as a data source.
Takeaway 6: Kaggle notebooks contain fewer error outputs than GitHub notebooks, but the crashes in Kaggle notebooks reflect challenges more specific to ML development, making them a more valuable resource in future research on the quality of ML notebooks.
5.3 Comparison with existing work
Frequent bugs in notebooks
De Santana et al. [8] find that environment and configuration issues are the most common bugs in Jupyter notebooks, often discussed on platforms like Stack Overflow. However, our study shows a higher prevalence of coding errors (52%), including API misuse, data confusion, and implementation issues. This difference likely reflects data sources: while the analysis from De Santana et al. focuses on user-reported issues, our study examines error outputs directly from executed notebooks. Our approach captures runtime failures, highlighting coding-related errors, whereas environment issues are often resolved before execution and thus less visible in our dataset.
ML/Python bugs
Our study finds that ML crashes are more frequent than general Python crashes (Kaggle: 71%, GitHub: 59%), in contrast with findings by Zhang et al. [11]. This discrepancy arises partly from differing definitions: they define ML bugs at a system level, encompassing all bugs in ML components of a system, while we focus specifically on bugs in code interacting with ML libraries. Additionally, both studies analyze error messages, but theirs are derived from job failures where code is tested, whereas ours come from cell outputs in notebooks, capturing issues during early development. This suggests that our study more directly reflects initial challenges faced by developers.
ML pipeline
Our study finds that the data preparation stage is the most crash-prone (GitHub: 29%, Kaggle: 26%) in the ML pipeline, similar to Islam et al. [7]. While their study includes bugs related to poor performance and incorrect outputs, not just crashes, and does not detail the impact of bugs within each stage, both studies highlight that data preparation remains the most challenging stage.
Root cause identification
Our study identifies new root causes of crashes in ML notebooks, particularly those related to notebook-specific issues like out-of-order execution and previous cell error. While out-of-order execution has been linked to reproducibility issues in prior work [2], it has not been explored as a crash root cause. Additionally, by analyzing crash information, context code, and cell outputs, we can distinguish crashes caused by data confusion from API misuse. Error messages and cell outputs reveal details about data shapes, types, and structures that reflect users’ understanding of their data, helping us identify crash origins. Examining cell outputs and documentation cells also enables us to identify intentional errors, such as API misuse shown in tutorial notebooks for instructional purposes.
6 Threats to validity
This section outlines potential threats to validity, addressing factors that may impact our findings.
6.1 External validity
Temporal limitation in GitHub dataset
The GitHub dataset includes notebooks from January 1, 2015, to March 31, 2022, which may not fully represent current notebook trends. Given the rapid evolution of coding practices and the tools used over time, notebooks from earlier years may reflect outdated coding styles or libraries. To mitigate this, we apply a filtering process to remove notebooks using older Python versions, improving the generalizability of the dataset.
Comparison of GitHub and Kaggle notebooks
The Kaggle dataset, which spans from January 1, 2023, to January 1, 2024, provides a more up-to-date snapshot of notebook trends. While the differing periods of GitHub and Kaggle datasets reduce redundancy in notebooks and associated crashes, the validity of comparison is threatened because ML practices may have changed. However, there is limited evidence of this effect because our results show no significant difference in exception types and root causes between Kaggle and GitHub.
KeyboardInterrupt exceptions
In our analysis, we exclude KeyboardInterrupt exceptions as these typically reflect user intervention rather than code issues. However, these errors could indicate infinite loops or unintended resource-heavy operations. By discarding these exceptions, we may overlook cases where such issues manifest.
6.2 Internal validity
Tutorial notebooks
The analyzed dataset contains various tutorial and educational notebooks. These tutorial notebooks often contain intentional crashes for didactic purposes. Such crashes could distort the crash patterns we aim to uncover. However, while these crashes are intentional, they still represent errors that users commonly believe occur in real-world scenarios, thus minimally affecting our results.
Manual labeling
We manually label ML notebook crashes, which can be prone to errors. To enhance the classification reliability, we use a structured two-stage process. In the labeling stage, three evaluators independently and iteratively categorize crashes, followed by discussions to synchronize findings, reduce bias, and promote consistency. In the reviewing stage, a reviewer validates the labels. Any disagreements are resolved through consensus.
Reproducibility of crashes
The error outputs we analyze represent real execution crashes but they are not necessarily reproducible due to factors like execution order, missing datasets, or environment replication [2]. While reproducing crashes could further validate our findings, it is challenging, especially for GitHub notebooks. Kaggle notebooks, with their accessible datasets and pre-configured environments, may enable reproducibility. In future work, we plan to reproduce crashes in Kaggle notebooks to validate our results and create a reproducible dataset for deeper analysis.
Clustering crashes
We cluster crashes with similar patterns based on their error values. We enhance this by removing specific substrings, like key names and numeric values. However, this results in an empty string in cases where error values only contain these specific substrings, which are consequently grouped into the same cluster. Examples are KeyError and SystemExit, which are very different errors. Although a theoretical threat, an investigation indicates that it only accounts for 0.38% of the dataset.
7 Conclusions and Future Work
In this paper, we present the first focused study on software crashes in ML notebooks from GitHub and Kaggle, uniquely identifying notebook-specific root causes, such as out-of-order execution and previous cell error. Our analysis classifies exception types, root causes, and ML pipeline stages, distinguishing between ML-specific and general Python bugs. Our findings reveal that the most common exceptions in ML notebooks are NameError, ValueError, and TypeError. The primary root causes are API misuse, data confusion, and notebook-specific issues (especially out-of-order execution). Notably, ML bugs account for more crashes than Python bugs, especially in Kaggle notebooks. Most crashes (58%) occur during the data preparation, model training, and evaluation/prediction stages of the ML pipeline. Additionally, GitHub and Kaggle exhibit different crash distributions across ML pipeline stages.
Future work will expand on these findings by constructing a reproducible dataset of ML crashes. Motivated by the results of this study, we also aim to develop an integrated, real-time bug detection tool tailored to ML development in the notebook environment, building on our previous work [39]. This tool will focus on ML bug detection while monitoring cell execution order and maintaining notebook semantics, helping developers catch errors as they code and enhancing productivity in ML development in notebooks.
Acknowledgments
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, and the Software Center Project 61.

### v2

@@
-Why do Machine Learning Notebooks Crash?
+Why do Machine Learning Notebooks Crash? An Empirical Study on Public Python Jupyter Notebooks
 Abstract
 Jupyter notebooks have become central in data science, integrating code, text and output in a flexible environment.
-With the rise of machine learning (ML), notebooks are increasingly used for ML prototyping and data analysis.
+With the rise of machine learning (ML), notebooks are increasingly used for prototyping and data analysis.
 However, due to their dependence on complex ML libraries and the flexible notebook semantics that allow cells to be run in any order, notebooks are susceptible to software bugs that may lead to program crashes.
-This paper presents a comprehensive empirical study focused on crashes in ML notebooks.
-We collect 64,031 ML notebooks containing 92,542 crashes from GitHub and Kaggle, and manually analyze a sample of 746 crashes across various aspects, including exception types and root causes.
-Our analysis highlights unique root causes related to notebook semantics, including out-of-order execution and previous cell error, that have not been thoroughly covered in earlier research.
-Furthermore, we categorize crashes as ML bugs or general Python bugs and examine how the crashes are distributed across different stages of the ML pipeline.
-Our analysis reveals that 87% of crashes are caused by API misuse, data confusion, notebook-specific issues, environment problems, and implementation errors.
-Crashes are more commonly related to ML bugs than general Python bugs, particularly in Kaggle notebooks, where over 70% of crashes are ML-related.
-The distribution of crashes across ML pipeline stages differs between the two platforms.
-Additionally, most crashes (58%) occur during data preparation, model training, and evaluation/prediction stages of the ML pipeline.
-GitHub and Kaggle exhibit different crash distributions across these stages.
+This paper presents a comprehensive empirical study focusing on crashes in publicly available Python ML notebooks.
+We collect 64,031 notebooks containing 92,542 crashes from GitHub and Kaggle, and manually analyze a sample of 746 crashes across various aspects, including crash types and root causes.
+Our analysis identifies unique ML-specific crash types, such as tensor shape mismatches and dataset value errors that violate API constraints.
+Additionally, we highlight unique root causes tied to notebook semantics, including out-of-order execution and residual errors from previous cells, which have been largely overlooked in prior research.
+Furthermore, we identify the most error-prone ML libraries, and analyze crash distribution across ML pipeline stages.
+We find that over 40% of crashes stem from API misuse and notebook-specific issues.
+Crashes frequently occur when using ML libraries like TensorFlow/Keras and Torch.
+Additionally, over 70% of the crashes occur during data preparation, model training, and evaluation or prediction stages of the ML pipeline, while data visualization errors tend to be unique to ML notebooks.
 Index Terms:
-Software bugs, machine learning, Jupyter notebooks, empirical study, crashes
+Software bugs, machine learning, Jupyter notebooks, python, empirical study, crashes
 1 Introduction
-Computational notebook environments like Jupyter notebooks have become indispensable in data science due to their interactive nature, offering real-time feedback during code execution [1].
-This interactivity makes them ideal for tasks such as data visualization, pattern exploration, and data analysis [2, 3].
-Notebooks also streamline documentation by combining text, code, execution results, and images into a single, cohesive environment, designed to enhance collaboration and reproducibility [2, 1].
+Jupyter notebooks, such as those used in Kaggle, JupyterLab, and Google Colab, have become indispensable in data science due to their interactive nature [1].
+They offer real-time feedback during code execution, making them ideal for tasks such as data visualization, pattern exploration, and data analysis [2, 3].
+Notebooks also streamline documentation by combining text, code, execution results, and visualizations into a single, cohesive environment, designed to enhance collaboration and reproducibility [2, 1].
 With the growing availability of data, machine learning (ML) and deep learning (DL) techniques have gained popularity for data modeling and decision-making within the broader field of data science [4, 5].
-In this context, ML notebooks have been widely adopted for ML prototyping and data analysis [6].
-Hence, ensuring the correctness of ML notebooks becomes essential to prevent software bugs from leaking into the later stages of ML development.
-ML notebooks are prone to software bugs with various underlying root causes.
-Such bugs commonly stem from the heavy reliance on ML libraries that are used for various stages in a ML pipeline including data processing, visualization, and building ML models or other data modeling algorithms [7].
-These powerful libraries often come with steep learning curves and varying abstraction levels, leading to errors in API usage, configuration, and compatibility.
-Furthermore, the flexible semantics of Jupyter notebooks, such as allowing code cells to be executed out of order or multiple times, could increase the likelihood of bugs [2, 1, 8].
+In this context, Jupyter notebooks, particularly Python-based ones, have been widely adopted for ML prototyping and data analysis [6].
+In industrial settings [7], Jupyter notebooks are often the starting point for developing ML applications before transitioning to production-ready ML pipelines, where much of the Python code used in notebooks is reused in large-scale distributed computing environments.
+This widespread adoption, even in mission-critical applications such as those at NASA Jet Propulsion Lab [8], underscores the importance of quality assurance for notebooks.
+Bugs that arise during the notebook stage may persist into later development stages, potentially affecting deployed products.
+Therefore, studying notebook bugs is essential to improve the overall reliability and robustness of ML-driven software.
+Software bugs in ML notebooks may be attributed to various underlying root causes.
+Such bugs commonly stem from the heavy reliance on ML libraries used for various stages in a ML pipeline including data processing and ML model building or other data modeling algorithms [9].
+These libraries often come with steep learning curves and varying abstraction levels, leading to errors in API usage, configuration, and compatibility.
+Furthermore, the flexible semantics of Jupyter notebooks, such as allowing code cells to be executed out of order, could increase the likelihood of bugs [2, 1, 10].
 Understanding these challenges is crucial for improving the debugging process and developing more robust tools for notebook environments.
-However, most previous studies [9, 7, 10, 11, 12] have been focused on ML scripts rather than ML programs developed in notebooks.
-Bugs in ML programs can manifest as low effectiveness, reduced efficiency, or crashes [9] which abruptly terminate the software.
+However, most previous studies [11, 9, 12, 13, 14] focus on ML scripts, which are standalone, linear code files (e.g., .py files), rather than ML programs developed in notebooks.
+In this paper, we refer to ML programs as a broader category that includes both ML scripts and ML notebooks.
+Bugs in ML programs can manifest in multiple ways, such as low effectiveness, reduced efficiency, or crashes [11] which abruptly terminate the execution.
 In Python, crashes occur as unhandled exceptions, representing the most severe bug symptom as execution is halted and needs immediate attention.
-Existing research [7, 12, 8] indicates that crashes are the most common symptoms of bugs in ML programs.
-Crashes are also more easily measurable due to accompanying error information that can be captured in the notebook environment.
-We conduct an empirical study to analyze 92,542 crashes in 64,031 ML notebooks by identifying their exception types, root causes, their association with ML libraries, and the ML pipeline stages where crashes typically occur.
-Further, we highlight root causes related to the improper use of notebook semantics (e.g., out-of-order execution) which has not been covered in previous research.
-Our goal is to enhance the use of Jupyter notebooks for ML development, guide the creation of better debugging tools, and improve the overall quality of ML notebooks.
+Existing research indicates that crashes are the most common symptoms of bugs in ML programs [9, 14, 10].
+Crashes are easier to investigate due to accompanying error information that can be captured in the notebook environment.
+We conduct an empirical study to analyze 92,542 crashes in 64,031 publicly available Jupyter notebooks for ML in Python (referred to as ML notebooks throughout the paper).
+We identify prevalent crash types and root causes, and investigate which ML libraries and pipeline stages most frequently contribute to crashes.
+Furthermore, we highlight root causes related to the improper use of notebook semantics (e.g., out-of-order execution) which has not been covered in prior research.
+Our long-term goal is to enhance the use of Jupyter notebooks for ML development, guide the creation of better debugging tools, and improve the overall quality of ML notebooks.
 To our best knowledge, this is the first study focused on analyzing crashes in ML notebooks.
-In particular, we examine crashes in notebooks collected from GitHub and Kaggle, addressing three main research questions.
-Additionally, we investigate how the characteristics of crashes in ML notebooks differ between these two platforms, providing a crosscutting aspect of our analysis.
-RQ1: What are the predominant exception types and root causes of crashes in ML notebooks?
+This study incorporates ML notebooks, used in Jupyter-based environments like Kaggle and JupyterLab, which share the same execution model and .ipynb format.
+We examine crashes in notebooks collected from representative repositories such as GitHub and Kaggle, which serve distinct user groups with different coding practices.
+Many data scientists and ML practitioners, particularly in exploratory and collaborative workflows, may lack strong software engineering skills and often commit notebooks with unresolved bugs [10].
+By focusing on those notebooks containing crash information, we gain insights into real-world debugging challenges and persistent errors in ML development, which are addressed along three main research questions.
+We also analyze how crash characteristics vary between GitHub and Kaggle, providing a crosscutting aspect of our analysis.
+RQ1: What are the predominant crash types and root causes of crashes in ML notebooks?
 Motivation: Crashes in notebooks often provide detailed error information including exception types that describe the direct trigger of the crash.
 The root cause, on the other hand, is the underlying reason behind a crash such as wrong usage of APIs, mismatched data expectations, or improper use of notebook semantics (e.g., out-of-order execution).
-Identifying the most common exception types, root causes, and their relationship help developers and researchers prioritize efforts on handling the most frequent issues in a systematic way.
+Identifying the most common crash types, root causes, and their relationship helps developers and researchers prioritize efforts to handle the most frequent issues in a systematic way.
 This makes the development of ML applications in notebooks more efficient and productive, while developers can still benefit from the flexibility of Jupyter notebooks.
-Results: The most common exception types in ML notebooks are NameError (local or global name not found), ValueError (function receives an argument of correct type but inappropriate value), and TypeError (operation or function is applied to an object with wrong type), attributing to 53% of crashes in total.
-The main root causes include API misuse (19%), notebook-specific issues (19%), data confusion (18%), environment problems (16%), and implementation errors (15%).
-Notebook-specific crashes, mostly due to out-of-order execution, are a leading cause on GitHub (20%) and third most common on Kaggle (18%).
-Our analysis of co-occurrences between exception types and root causes reveal that NameError is primarily (56%) caused by notebook-specific factors.
-API misuse and data confusion are the primary causes of a wide range of exception types.
-RQ2: To what extent are crashes in ML notebooks caused by ML bugs compared with general Python bugs?
+Results: In our study, the most common crash types in ML notebooks are Variable Not Found (variable name not found), Invalid Argument (function receives an inappropriate argument), and IO Error (errors when handling files), attributing to 37.3% of crashes in total.
+We identify ML-specific crash types including Tensor Shape Mismatch, Data Value Violation, Out of Memory, Unsupported Broadcast, Feature Name Mismatch, and Model Initialization Error.
+The main root causes include API misuse (20.9%) and notebook-specific issues (19.4%, mostly due to out-of-order execution).
+Our analysis reveals that name-related errors (i.e., undefined names) are primarily caused by notebook-specific issues.
+API misuse, implementation error, and data confusion are the primary causes of a wide range of crash types.
+RQ2: To what extent are crashes in ML notebooks caused by ML bugs and by which ML libraries?
 Motivation: Crashes in ML notebooks can result from ML-related bugs and general Python bugs.
-We examine to what extent these crashes are caused by ML bugs, including those directly caused by the wrong use of APIs in ML libraries and other errors that arise from the code using ML APIs or variables.
+We examine to what extent these crashes are caused by ML bugs, including those that arise from code using ML APIs or involving variables, objects, or other components derived from ML APIs.
+Additionally, we determine whether a crash results from the direct use of an ML library to identify which libraries are more often associated to crashes and pose greater debugging challenges.
 This helps guide better ML API usage practices and improves ML library stability.
-Results: ML bugs account for 64% of ML notebook crashes, with 59% on GitHub and a significantly higher proportion (71%) on Kaggle.
-The primary causes of ML-related crashes are API misuse (23%), data confusion (23%), and notebook-specific issues (21%), while general Python bugs are mostly due to environment (24%) and implementation (24%) errors.
+Results: In our study, ML bugs account for 67.7% of ML notebook crashes, with 62.8% on GitHub and a significantly higher proportion on Kaggle (73.0%).
+The primary causes of ML-related crashes are API misuse (25%), data confusion (20%), and notebook-specific issues (20%), while general Python bugs are mostly due to implementation (29%) and environment (28%) errors.
 API misuse is more common in ML-related crashes compared to Python crashes, highlighting the increased difficulty of using ML libraries correctly.
-RQ3: In which stages of the ML pipeline are crashes in notebooks most likely to occur?
+In particular, tensorflow/keras and torch are shown to be the most common source of crashes with the highest crash ratio relative to its usage.
+RQ3: In which stages of the ML pipeline are crashes in notebooks most frequently to occur?
 Motivation: Studying which stages of the ML pipeline are most prone to crashes, such as data preparation, model construction, or model training stages, helps developers and researchers focus on improving robustness in critical stages, and optimize development resources accordingly.
-Results: Crashes predominantly occur during data preparation (27%), model training (16%), and evaluation/prediction (15%) stages.
-GitHub experiences 3.5 times as many crashes unrelated to any ML pipeline stages compared to Kaggle.
-The two platforms also exhibit distinct crash distributions across pipeline stages.
+This does not imply that all ML notebooks follow a complete pipeline.
+For example, some may exclusively focus on data preparation, omitting other stages such as model construction or training.
+Therefore, we analyze which stages of a general ML pipeline caused the crash, regardless of whether a notebook implements all of them.
+Results: In our study, crashes predominantly occur during data preparation (33.0%), model training (19.4%), and evaluation/prediction (18.4%) stages.
+GitHub has more crashes during environment setup and data preparation, whereas more crashes are observed during the evaluation/prediction stage in Kaggle.
 Additionally, over 70% of the crashes during model training and evaluation/prediction are caused by ML bugs.
-In data preparation, Kaggle has a higher proportion of ML-related crashes compared to GitHub.
-In contrast, in the environment setup stage, GitHub notebooks show more ML-related crashes.
-Replication package. All artefacts can be found at https://github.com/PELAB-LiU/jupyter_nbs_empirical.
-Organization.
-The remaining of this paper is organized as follows.
-Section 2 provides background information and discusses related work relevant to our study.
-Section 3 outlines the experimental setup, detailing the methods used for data collection, processing, and analysis.
-Next, Section 4 presents the results of our study.
-Section 5 discusses our main findings and addresses their implications and significance.
-Finally, Section 6 discusses the potential threats to validity of this study and Section 7 concludes the paper.
+Replication package. All experimentation artifacts can be found on GitHub [15], and data on Zenodo [16].
 2 Background and related work
@@
 Study Analysis dimensions Targets Libraries Data sources
-Zhang et al. [9] Root cause, symptom ML/DL scripts TensorFlow SO posts, GH issues
-Islam et al. [7] Bug type, root cause, symptom, ML pipeline ML/DL scripts Caffe, Keras, TensorFlow, Theano, PyTorch SO posts, GH commits
-Humbatova et al. [10] Fault taxonomy (root cause) ML/DL scripts TensorFlow, Keras, PyTorch SO posts, GH commits/issues, interviews, survey
-Zhang et al. [11] Failure type, root cause ML/DL scripts Not limited Failure messages from a Microsoft DL platform, interviews
-Morovati et al. [12] ML/non-ML bugs, root cause and symptom of ML bugs ML/DL systems TensorFlow, Keras, PyTorch GH issues
+Zhang et al. [11] Root cause, symptom DL scripts TensorFlow SO posts, GH issues
+Islam et al. [9] Bug type, root cause, symptom, ML pipeline DL scripts Caffe, Keras, TensorFlow, Theano, PyTorch SO posts, GH commits
+Humbatova et al. [12] Fault taxonomy (root cause) DL scripts TensorFlow, Keras, PyTorch SO posts, GH commits/issues, interviews, survey
+Zhang et al. [13] Failure type, root cause, ML pipeline DL scripts Not limited Failure messages from a Microsoft DL platform, interviews
+Morovati et al. [14] ML/non-ML bugs, root cause and symptom of ML bugs ML(DL) systems TensorFlow, Keras, PyTorch GH issues
 Ahmed et al. [17] Bug type, root cause, symptom Data analysis programs Most tagged ones on SO SO posts, GH commits/issues
@@
 2.1 Terminology and motivating example
-First we clarify some key terms following the definitions provided in a previous study [18].
+First, we clarify some key terms following the definitions provided in a previous study [22].
 We use bugs or faults as inherent defects that result in incorrect or unexpected behaviors, known as errors.
 If left unaddressed, these errors can lead to failures or crashes.
-In Python, crashes are represented by exceptions, typically accompanied by error information.
+In Python, crashes manifest as exceptions, typically accompanied by error information.
 Jupyter notebooks are the most common medium of ML prototyping and data analysis [6].
 Notebooks are more flexible compared to, for example, Python scripts by offering additional interactive features.
-Figure 1 shows an example excerpt of a ML notebook that trains and evaluates a neural network.
-A notebook typically consists of multiple cells, which can either contain documentation or executable code.
+Figure 1 shows an example excerpt of an ML notebook that trains and evaluates a neural network.
+A notebook typically consists of multiple cells, which can contain either documentation or executable code.
 When a user runs a code cell, Jupyter uses a kernel to execute it.
@@
 While this supports iterative and exploratory development, it can lead to unexpected behavior if cells are executed out of sequence.
-For example, the notebook in Figure 1 can be executed in reverse, such that the code that evaluates the model is executed before the code that trains the model.
+For example, the notebook in Figure 1 can be executed in reverse, such that the code that evaluates the model is executed before the code that compiles and trains the model.
+In our example, this would cause a crash.
 2.2 Studies of bugs in ML or data science programs
@@
 Moreover, no studies focus on program crashes.
-Our study addresses these gaps by analyzing crashes in ML notebooks with investigation of how notebook semantics (e.g., out-of-order execution, or previous cell error) impact crashes.
-Although prior work [2] has linked out-of-order execution to reproducibility issues, we are the first to identify various notebook-specific semantic issues as root causes of crashes.
+Our study addresses these gaps by analyzing crashes in Python ML notebooks with investigation of how notebook semantics (e.g., out-of-order execution, or previous cell error) impact crashes.
+Although prior work [2] has linked out-of-order execution to reproducibility issues, we are the first to identify notebook-specific semantic issues as root causes of crashes.
 By examining error information in cell outputs, we distinguish ML bugs from general Python bugs and pinpoint the stages of the ML pipeline where these crashes occur.
@@
 This allows us to capture bugs as they occur, providing more accurate insights into the nature of crashes.
-By analyzing these crash information, we can identify and diagnose issues more effectively.
-Furthermore, many previous studies have focused on bugs in ML/DL programs that use specific ML libraries [9, 7, 10, 13, 12], or defined ML bugs as all errors that occur in ML components in a system [12].
-In contrast, our approach is not restricted to a few ML libraries.
-Instead, we examine a broad range of code from our dataset and identify popular ML libraries.
-We define ML bugs as bugs in the crashed code that interacts with such libraries, which provides a more direct and precise link between crashes and ML libraries.
+By analyzing this crash information, we can identify and diagnose issues more effectively.
+Furthermore, many previous studies focused on bugs in ML/DL programs that use specific ML libraries [11, 9, 12, 17, 14], or defined ML bugs as all errors that occur in ML components in a system [14].
+In contrast, we examine a broad range of code from our dataset and identify popular ML libraries.
+As we investigate ML bugs which cause crashes of code that interacts with such libraries, our approach provides a more direct link between crashes and ML libraries.
 Figure 2: Overview of the data collection, filtering, sampling, and manual analysis for Kaggle and GitHub notebooks.
@@
 This section outlines the experimental setup of our study, as shown in Figure 2.
-We begin with the data sources and collection methods, followed by the filtering criteria for identifying the relevant population.
+We begin with the data collection methods, followed by the filtering criteria for identifying the relevant population.
 Next, we describe the sampling strategy used to select a representative sample.
-Lastly, we describe the process for manual analysis.
+Lastly, we describe the process for manual analysis and resampling.
 3.1 Data sources and mining method
@@
 Python notebooks
-This study focuses on notebooks that use Python.
-To identify the programming language of each notebook, we use the Guesslang library [^1] https://github.com/yoeo/guesslang used in previous works [22, 23] for similar filtering steps.
+This study focuses on Python notebooks.
+To identify the programming language of each notebook, we use the Guesslang library [26] used in previous works [27, 28] for similar filtering steps.
 The top three most commonly used languages in GitHub notebooks are Python (95.35%), Julia (1.22%), and R (0.92%).
@@
 Hence, we parse and filter all notebooks to retain only those that contain crashes in their cell outputs.
-Specifically, we filter notebooks containing at least one cell such that 1) its cell_type is code, 2) it comes with output whose output_type is error.
+Specifically, we filter notebooks containing at least one cell such that 1) its cell_type is code, 2) its output_type is error.
 After applying this filter, we observe that 10.67% of the collected GitHub notebooks and only 2.61% of Kaggle notebooks contain crashes.
@@
 Valid notebooks
-We exclude notebooks with syntactically incorrect code because such errors prevent the code from running at all, making them irrelevant to the actual crashes developers often face.
-We focus on crashes that occur during code execution, which reveal deeper issues affecting program quality.
+As our study targets runtime crashes during execution, we exclude notebooks with syntactically incorrect code.
+Syntax errors prevent execution, making the notebook non-reproducible and unreliable (garbage-in, garbage-out).
+Furthermore, we investigate notebook-specific issues, such as previous cell errors, focusing on previous cells that execute incorrectly rather than contain syntactically invalid code.
+Including crashes caused by invalid code would highlight problems that can be resolved using basic static analysis tools, thus yielding no insights into the unique challenges faced during ML notebook development.
+Additionally, notebooks with persistent syntax errors are typically of lower quality, often containing incomplete work, and may not reflect typical ML development workflows.
 To check for syntactic validity, we use Python 3 Abstract Syntax Tree (AST) module, which parses code according to Python 3 grammar.
@@
 ML notebooks
-Jupyter notebooks have various uses.
+Jupyter notebooks can be used for various applications.
 This paper focuses specifically on ML notebooks, defined as those that use at least one ML library.
@@
 We extract and rank all libraries imported by Kaggle notebooks based on frequency.
-The top 50 libraries, used by $\sim$95of all Kaggle notebooks, are reviewed by two authors to determine their relevance to ML.
-From this, we identify 24 popular ML libraries (see Table II), which we use to identify ML notebooks in our dataset.
-TABLE II: The 24 most popular ML libraries identified from 166,655 Kaggle notebooks together with their usage frequencies (%), grouped by their main functionality.
+The top 50 libraries, used in $\sim$95of all Kaggle notebooks, are independently reviewed by two authors to determine their relevance to ML.
+Full agreement is reached in identifying 24 popular ML libraries (see Table II).
+We then use these libraries to identify ML notebooks in our dataset.
+TABLE II: Top 24 popular ML libraries identified from 166,655 Kaggle notebooks with their usage frequencies, grouped by their main functionalities. DL libraries are marked in bold.
 Theme ML libraries
@@
 Normally, error values of crashes can be extracted from notebooks without extra processing.
-But, in some cases, developers substitute error values with non-descriptive messages like ignored or null.
+But, in some cases, developers substitute error values with non-descriptive messages like “ignored” or “null”.
 We replace them with a more descriptive message extracted from the traceback of the crash.
@@
 Our study requires a manual labeling and review process to analyze ML notebook crashes.
-Since investigating the entire population is infeasible, we select a meaningful sample of crashes for manual analysis similarly to De Santana et al.
-[8].
+Since investigating the entire population is infeasible, we select a meaningful sample of crashes for manual analysis.
 We apply proportional stratified sampling [29], a strategy that builds on top of simple random sampling.
@@
 An example of clustering results using Jaccard similarity is presented in Table III.
-We chose a threshold of 0.7 to perform Jaccard similarity based on preliminary experiments.
+We choose a 0.7 threshold for Jaccard similarity, based on its use in a previous study [32] in the context of code clone detection.
+We also test smaller and larger thresholds (ranging from 0.5 to 0.9 in 0.1 intervals).
+After evaluating 50 clusters formed at each threshold, we find that the smaller thresholds lead to clusters with mixed error messages, while the larger ones create too many similar clusters.
+The 0.7 threshold provides a coherent and meaningful grouping, making it the optimal choice for our analysis.
 We first cluster all crashes (both GitHub and Kaggle) to get a comprehensive view of common patterns, then calculate clusters for GitHub and Kaggle datasets separately to identify platform-specific differences in crash behaviors.
@@
 Separate statistics for each platform reveal 7,222 clusters for GitHub and 640 for Kaggle.
-TABLE III: Jaccard similarity result of two error values.
-Error value (original) Error value (processed) Jaccard similarity
-name ’x’ is not defined name is not defined 1.0 (the same)
-name ’head’ is not defined name is not defined
+TABLE III: Jaccard similarity result (JS) of two error values.
+Error value (original) Error value to compare JS
+Input 0 of layer ”sequential_1” is incompatible with the layer: expected shape=(None, 32), found shape=(None, 38) input of layer is incompatible with the expected shape none found 0.79
+Input 0 of layer sequential_1 is incompatible with the layer: expected ndim=3, found ndim=2.
+Full shape received: (None, None) input of layer is incompatible with the expected ndim found full shape received none
+TABLE IV: Different clusters within Value Error exception.
+Cluster Example error value Crash type
+2255 operands could not be broadcast together with shapes (25,200) (25,50) Unsupported broadcast
+559 Classification metrics can’t handle a mix of binary and continuous targets Data value violation
+720 Input 0 of layer ”sequential” is incompatible with the layer: expected shape=(None, 4), found shape=(None, 128) Tensor shape mismatch
+6862 cannot reshape array of size 0 into shape (2,2)
+6473 shapes (25,) and (26,1) not aligned: 25 (dim 0) != 26 (dim 0)
+TABLE V: Classification dimensions of crashes in ML notebooks.
+RQ Dimension Category Description
+RQ1 Crash type Variable Not Found (VNF) A referenced variable is not defined or accessible. A refined type of Name Error.
+Invalid Argument (IARG) A function receives an argument with an incorrect type, format, or value, or has missing or misordered arguments, violating API constraints.
+IO Error (IO) A crash occurs when reading from or writing to a file, e.g., file/path not found, no permission to the file, invalid file type, or unable to parse or decode the file.
+Module Not Found (MODULE) A required library or module is missing or not installed. A refined type of Name Error.
+Tensor Shape Mismatch (TSHAPE) Tensor shapes are incompatible for the intended operation. A refined type of Value Error.
+Data Value Violation (DVIOL) Dataset contains unexpected or invalid values that violate API constraints or user assumptions. A refined type of Value Error.
+Out of Memory (OOM) The available memory of GPU/CPU is insufficient to execute the operation.
+Environment Error (ENV) Issues related to system setup or import failures.
+Request Error (RERR) A failure occurs when making an external API or network request.
+Unsupported Broadcast (BCAST) Broadcasting fails due to incompatible array or tensor shapes. A refined type of Value Error.
+Feature Name Mismatch (FNAME) Feature names do not match expected names or formats in ML model structure or dataset. A refined type of Value Error.
+Model Initialization Error (INIT) Crash during model creation due to incorrect parameters or architecture.
+Python built-in exception types (e.g., Attribute Error(ATTR), Key Error(KEY), Value Error(VALUE), Type Error(TYPE), Name Error(NAME), Index Error(INDEX), Runtime Error(RUNTIME), System Error, Zero Division Error) These are exceptions that occur during the execution of Python code, typically caused by issues like attribute access errors, illegal dictionary keys, invalid values, type mismatches, undefined function/class names, out-of-bounds indexing, execution failures, system-related issues, division by zero, and so on.
+More details can be found in the Python documentation.
+Root cause API misuse (API) Incorrect usage of an API, e.g., passing invalid arguments or calling functions improperly.
+Notebook-specific issue (NB) Errors that stem from the unique execution behavior of notebooks, e.g., out-of-order execution, and unresolved errors in prior cells (previous cell error).
+Data confusion (DATA) Misinterpretation of data shapes, structure, or values, leading to unexpected errors.
+Implementation error (IMPL) Mistakes in code logic, e.g., incorrect variable usage, function calls, or algorithm design.
+Library issue (LIB) Errors originating from external libraries, including bugs or deprecated APIs.
+Environment issue (ENV) Issues due to incorrect environment setting, e.g., missing/incompatible dependencies, file path errors, or external control.
+ML model Mismatch (MODEL) A mismatch between a saved ML model and the expected model definition in the code.
+Insufficient resource (RSC) Crashes due to resource constraints, e.g., running out of memory or disk space.
+Unknown (UNK) Errors for which the root cause cannot be determined.
+RQ2 ML/Python bug ML bug Errors in code using ML libraries or related components.
+Python bug Errors from general Python code, unrelated to ML libraries.
+Library cause Specific ML libraries The crash is caused by using one/more ML libraries (see Table II).
+None The crash is neither ML-related bug nor caused by directly using any ML libraries.
+RQ3 ML pipeline Environment setup (ENVS) Errors related to environment setup, e.g., missing modules or configuration issues.
+Data preparation (DATAP) Errors occurring during data preprocessing or preparation for analysis and training.
+Data visualization (DATAV) Errors encountered while visualizing data.
+Model construction (MCONS) Issues related to building, compiling, or visualizing ML models.
+Model training (TRAIN) Errors that arise during model training, including hyperparameter tuning and validation.
+Evaluation/ prediction (EVAL) Errors related to evaluating model performance or making predictions.
 We further visualize the clusters in relation to exception types.
-We observe that multiple exception types have similar error values as Table IV shows.
+We observe that one exception type may have different patterns that lead to different crash types as Table IV shows.
 This is likely because Python exceptions can be customized and raised by users or library developers under various names or reasons.
-As a result, exceptions with different names may represent similar issues.
 This finding motivates our approach of clustering based on error values, rather than using exception types directly.
@@
 This results in 66 clusters for GitHub and 51 for Kaggle overall.
-TABLE IV: Top 5 clusters across GitHub and Kaggle datasets.
-The first column shows cluster sizes (CS) with the number of exception types (#ET) per cluster.
-The second column lists the top 3 exception types and an example error value.
-Python built-in exception types are explained, while customized exceptions are marked as “self-defined.”
-CS (#ET) Top 3 ET (exception types) Example error value
-25070 (3) NameError(25050): local/global name not found UndefinedVariableError(19): self-defined RemoteError(1): self-defined
-name ’sc’ is not defined
-6107 (6) AttributeError(6072): attribute reference/assignment fails ModuleAttributeError(18): self-defined TypeError(13): operation/function apply to wrong type object
-’numpy.ndarray’ object has no attribute ’sample’
-5136 (3) ModuleNotFoundError(4568): module could not be located ImportError(567): cannot load a module DependencyNotinstalled(1): self-defined
-No module named ’src.utils.compute_overlap’
-4620 (83) KeyError(3911): a mapping key is not found SystemExit(349): raised by the sys.exit() ValueError(48): argument with right type but wrong value
-’Hero Over’
-3410 (16) FileNotFoundError(3230): file/directory not exist IOError(79): aliases of OSError OSError(28): system-related error,  e.g., ’disk full’
-No such file or directory: ’GloVe/glove.840B.300d.txt’
-TABLE V: Classification dimensions of crashes in ML notebooks.
-RQ Dimension Category Description
-RQ1 exception type Python built-in exception types (e.g., NameError, ValueError, TypeError, KeyError) These are exceptions that occur during the execution of Python code, typically caused by issues like undefined variables, incorrect data types, or illegal key errors.
-Some examples can be seen in Table IV, more details can be found in the Python documentation.
-others (e.g., APIArgumentError, EnvironmentError, ResourceError) These consist of library customized exceptions.
-For example, APIArgumentError arise when argument values fail to meet expected constraints of an API, EnvironmentError include import exceptions and system configuration issues, ResourceError contain exceptions such as out of memory or out of disk space.
-root cause data confusion (DATA) Errors occur when the expected data structure (e.g., shapes, value types, feature names) does not match the actual input, or previously saved ML models (considered as external data) fails to align with the model defined in the code (e.g., model structure).
-API misuse (API) Using an API incorrectly, e.g., providing wrong arguments or applying an API improperly.
-notebook-specific issues (NB) Errors that stem from the unique execution behavior of notebooks, e.g., out-of-order execution, and unresolved errors in prior cells (previous cell error).
-implementation error (IMPL) Faults in the coding logic, e.g., incorrect implementation or missing appropriate conditions.
-library change (LIB) Errors caused by updates to an API or libraries, making current code incompatible.
-environment issues (ENV) Issues due to incorrect environment setting, e.g., missing/incompatible libraries, file path errors.
-intentional (INTENT) Errors deliberately introduced by users, often for testing or education purposes.
-insufficient resources (RSC) Crashes due to resource constraints, like running out of memory or disk space.
-unknown (UNK) Errors for which the root cause cannot be determined.
-RQ2 ML/Python bug ML bug Errors in code using ML libraries or related components.
-Python bug Errors from general Python code, unrelated to ML libraries.
-RQ3 ML pipeline environment setup (ENVS) Errors occurring when setting up the environment, e.g., missing modules or configuration issues.
-data preparation (DATAP) Errors occurring during data preprocessing or preparation for analysis and training.
-data visualization (DATAV) Errors encountered while visualizing data.
-model construction (MCONS) Issues related to building, compiling, or visualizing ML models.
-model training (MTRAIN) Errors that arise during model training, including hyperparameter tuning and validation.
-evaluation/ prediction (EVAL) Errors related to evaluating model performance or making predictions.
-no ML pipeline (NONE) Errors unrelated to ML pipeline (e.g., notebooks for tutorials, education, or simulations).
 Proportional-to-cluster-size sampling
@@
 We calculate sample sizes and analyze sampled crashes separately for GitHub and Kaggle because they may exhibit different behaviors.
-The desired sample sizes are determined with a 95% confidence interval (CI) and a 5% margin of error (MoE), aiming for results that are 95% accurate within a 5% margin of error from the entire population.
+The desired sample sizes are determined with a 95% confidence interval (CI) and a 5% margin of error (MoE).
 This results in sample sizes of 360 for GitHub and 332 for Kaggle crashes.
+The details of the sample size calculation process can be found in our GitHub repository [15].
 We then use these sample sizes and the identified clusters (GitHub: 66, Kaggle: 51) to conduct proportional sampling based on cluster size.
@@
 Table V shows the 4 dimensions considered in our analysis and how they relate to our research questions.
-EXCEPTION TYPE
+CRASH TYPE
 The first dimension categorizes crashes based on the exceptions that trigger them, originally shown as the error names in the notebook outputs (see Figure 1).
@@
 We manually refine the raw exception types for two reasons.
-First, we distinguish between Python’s built-in exceptions and those defined by third-party libraries.
-Our analysis reveals that 7.21% of crashes on GitHub and 7.31% on Kaggle have names differing from Python’s built-in types, indicating these exceptions are likely created by third-party libraries or users.
-By refining exception types, we can better identify library-specific exceptions.
-Moreover, as shown in Table IV, different exception types may exhibit similar patterns, while the same exception type may represent different triggers.
+First, we distinguish between Python’s built-in exceptions and those defined by library developers or users.
+Our analysis reveals that 7.21% of crashes on GitHub and 7.31% on Kaggle have names differing from Python’s built-in types.
+By refining exception types, we can better identify customized exceptions.
+Moreover, as shown in Table IV, one exception type may exhibit different patterns.
 Users can define and raise exception types freely, which may not always accurately represent the true trigger of a crash.
@@
 The flexibility of notebooks, which allows cells to be executed at any time, can introduce further complications.
-We base our initial root cause categories on prior studies [7, 11, 8].
-We also explore how notebook semantics, such as out-of-order execution and unresolved errors in previous cells, contribute to crashes.
-By categorizing crashes by root causes, we capture a broad range of crash sources, from misunderstanding of input data or APIs to environment-specific and notebook-specific issues, obtaining a comprehensive understanding of why the crashes happen.
+We base our initial root cause categories on prior studies (API misuse [9, 13], data confusion [13], implementation error [13, 10], library issues [9], environment issue [10], insufficient resource [10]).
+We extend this initial set of categories based on further insights from our pre-study.
+We discover how notebook semantics, such as out-of-order execution and previous unresolved errors, contribute to crashes.
+Additionally, we identify ML model mismatches as a root cause, occurring when a previously saved model is incompatible with the ML model defined in the current code.
+By categorizing crashes by root causes, we capture a broad range of crash sources, from misunderstanding of data or APIs to environment and notebook-specific issues, obtaining a comprehensive understanding of why the crashes happen.
 ML/PYTHON BUG
@@
 [14] compare ML bugs and non-ML bugs by identifying ML bugs as all bugs within ML components of DL systems.
-In contrast, our study focuses on the ML components themselves, aiming to evaluate how often crashes stem from direct interactions with ML libraries.
+In contrast, our study focuses on the ML components themselves, aiming to evaluate how often crashes stem from interactions with ML libraries.
 We define ML bugs as crashes that occur in code using ML libraries or objects derived from them.
 By distinguishing ML bugs and general Python bugs, we can better understand the challenges unique to ML development, including the use of ML APIs, model handling, and data processing tasks.
+LIBRARY CAUSE
+This dimension labels which ML library is directly responsible for the crash.
+In our ML/Python bug classification, a crash is considered an ML bug if the code involves any ML APIs or components, regardless of whether the issue directly stems from ML libraries.
+This dimension focuses specifically on crashes caused by ML libraries, meaning that the crash must occur as a direct use of the library.
+As such, this is a subset of ML bugs.
+This dimension can contribute to determine which ML libraries are more challenging to use.
 ML PIPELINE
 In this dimension, we categorize crashes based on where they occur during the ML pipeline to identify critical stages.
-Building on the framework from Islam et al.
-[7] and a pilot study of our dataset, we define key stages including environment setup, data preparation, data visualization, model construction, model training, and evaluation/prediction.
-This dimension is crucial for analyzing where in the ML pipeline crashes are more frequent and potentially harder to resolve.
+We define key stages building on the framework from a prior study by Zhang et al.
+[13] (environment setup, data preparation, training, and evaluation/prediction stages) and a pre-study of our dataset (data visualization and model construction stages).
+This categorization does not imply that all notebooks adhere to a strict pipeline.
+For example, some only implement a subset of stages, while others contain the same stage multiple times (e.g., when they try and compare different preprocessing methods).
+Rather, it provides a structured way to analyze where crashes occur within a general ML workflow and where crashes in the ML pipeline are more frequent and potentially harder to resolve.
 3.5 Manual labeling process
@@
 We adopt grounded theory methods [35], incorporating first-cycle and second-cycle coding phases [36].
-We begin with an initial set of categories for each dimension, derived from prior studies [11, 7, 12] and our pilot study, and iteratively refines them as needed to maintain classification accuracy.
+We first label crashes based on predefined initial categories for each dimension (see Section 3.4).
+We then iteratively refine these categories whenever new crash patterns emerge that do not fit within the existing ones, particularly for Crash type and Root cause.
+We do this by adding new categories and updating existing ones, which is common in grounded theory [35].
+For example, the crash type Tensor Shape Mismatch was originally included in Value Error and later split into a separate category.
+This creates more consistent and well-separated categories.
 During the first-cycle coding, three evaluators independently label an initial set of 50 crashes from the sampled dataset.
@@
 To further improve consistency, we organize crashes from both GitHub and Kaggle datasets by clusters, which groups similar crashes together for labeling and review.
-Each crash is categorized based on the predefined dimensions, with new categories introduced only when necessary.
 Additionally, evaluators reference all notebooks containing the sampled crashes to access relevant crash contexts.
@@
 During the second-cycle coding, the remaining crashes are labeled independently.
-Regular meetings are held to discuss and reach consensus on any disagreements, promoting consistent standards across the entire dataset.
-In the reviewing stage, a separate evaluator conducts reviewing to validate the labeling results.
+Regular meetings are held to discuss and reach a consensus on any disagreements, promoting consistent standards across the entire dataset.
+In the reviewing stage, a separate evaluator validates the labeling results.
 In cases of disagreement, discussions take place to resolve conflicts, and if needed, all three evaluators come together to reach a final decision.
+TABLE VI: Changes in the number and percentage of labeled crashes after reviewing the final of sampled crashes (746 crashes after resampling).
+RQ Dimension Number Percentage
+RQ1 Crash type 31 4.16%
+Root cause 72 9.65%
+RQ2 ML/Python bug 24 3.22%
+Library cause 32 4.29%
+RQ3 ML pipeline 64 8.58%
+3.6 Resampling
+After manually labeling and reviewing the data, we identify 123 cases in which the crashes do not fit into any stage of the ML pipeline.
+The majority of these cases originate from tutorial notebooks, educational materials, or physics simulations.
+Additionally, we find four irrelevant crashes (e.g., cells containing random text, exceptions raised to print comments).
+As these instances do not represent actual crashes in ML development, we exclude them from our analysis.
+To maintain statistically significant sample sizes, we use a resampling approach similar to proportional-to-cluster-size sampling (Section 3.3), where replacements are drawn from the same clusters where exclusions occur.
+We then verify whether the replacements should be excluded as well.
+If such cases exist, we iteratively perform resampling until we reach the target sample sizes (GitHub: 390, Kaggle: 356).
+Afterwards, we manually label the 127 resampled crashes following the same process (see Section 3.5).
 The change rates before and after reviewing are shown in Table VI.
-The process sometimes results in updates to earlier categories, particularly as the classification standards are refined during review.
+The reviewing process sometimes results in updates to earlier categories.
 For example, many changes are made to categorize crashes in ML pipeline based on the stage where the code is written, rather than by its functionality, which leads to a high rate of reclassification.
 On the other hand, the high change rate of Root cause reflects the challenges of accurately identifying the root causes of the crashes.
-TABLE VI: Changes of labeled crashes after reviewing.
-RQ Dimension Changed total Percentage
-RQ1 exception type 37 14.51%
-root cause 81 31.76%
-RQ2 ML/Python bug 25 9.80%
-RQ3 ML pipeline 82 32.16%
-After manually labeling and reviewing, we find four irrelevant crashes, such as cells containing random text like “end,” or exceptions raised to print comments like “Done!” These do not represent actual crashes and are excluded from our analysis, resulting in 389 labeled crashes from GitHub and 353 from Kaggle.
-Despite this reduction, the sample sizes remain statistically significant with 95% confident level.
 4 Results
@@
 For distributions with small sample sizes ($<5$), we apply Barnard’s exact test [37], using Bonferroni correction [38] when the categories are non-binary.
-4.1 RQ1: What are the predominant exception types and root causes of crashes in ML notebooks?
-(a) Exception types
+When specifying our results, we refer to GitHub as GH and Kaggle as KG.
+4.1 RQ1: What are the predominant crash types and root causes of crashes in ML notebooks?
+(a) Crash types (see Table V for the full names)
 (b) Root causes (see Table V for the full names)
@@
 Crashes in ML notebooks stem from various types of exceptions.
-We refine and enhance the raw exception types typically shown in the crash information.
-Additionally, we complement these classes by identifying their underlying root causes and discussing how they relate to exception types.
+We refine and enhance the raw exception types typically shown in the crash information and form crash types.
+We then complement these classes by identifying their underlying root causes and discussing their relations.
 The results are visualized in Figure 3 and Figure 4.
-Exception types
-3(a) highlights the top 10 most frequent exception types.
-The most common exception type is NameError, accounting for 30% of cases (GitHub: 31%, Kaggle: 28%), followed by ValueError at 12% (GitHub: 11%, Kaggle: 12%), TypeError at 11% (GitHub: 10%, Kaggle: 12%), IOError at 9% (GitHub: 8%, Kaggle: 11%), and APIArgumentError also at 9% (GitHub: 10%, Kaggle: 8%).
-The $\chi^{2}$ test shows no significant difference in exception type distributions between GitHub and Kaggle ($p>0.05$).
-Finding 1.1: The most common exception types of ML notebook crashes are NameError (30%), ValueError (12%), and TypeError (11%).
+Crash types
+3(a) highlights the most frequent crash types in both GitHub and Kaggle and crash types that are unique to ML.
+The most common crash type is Variable Not Found, accounting for 17.8% of cases (GH: 18.5%, KG: 17.1%), followed by Invalid argument at 9.8% (GH: 10.3%, KG: 9.3%), IO Error at 9.7% (GH: 9.0%, KG: 10.4%), Module Not Found at 8.0% (GH: 9.7%, KG: 6.2%), and Attribute Error at 7.0% (GH: 7.4%, KG: 6.5%).
+Finding 1.1: The most common crash types of ML notebook crashes are Variable Not Found (17.8%), Invalid argument (9.8%), and IO Error (9.7%).
+Moreover, crash types Tensor Shape Mismatch (5.5%; GH: 4.6%, KG: 6.5%), Data Value Violation (3.5%; GH: 3.9%, KG: 3.1%), Out of Memory (2.1%; GH: 1.3%, KG: 3.1%), Unsupported Broadcast (1.6%; GH: 1.3%, KG: 2.0%), Feature Name Mismatch (1.2%; GH: 1.5%, KG: 0.8%), and Model Initialization Error (0.7%; GH: 0.5%, KG: 0.8%) are ML-specific crash types.
+The $\chi^{2}$ test shows no significant difference in crash type distributions between GitHub and Kaggle ($p>0.05$).
+Finding 1.2: ML-specific crashes types (14.6%) include Tensor Shape Mismatch, Data Value Violation, Out of Memory, Unsupported Broadcast, Feature Name Mismatch, and Model Initialization Error.
 Root causes
-3(b) visualizes the distribution of root causes.
-The most common ones are API misuse (19% overall; GitHub: 18%, Kaggle: 21%), notebook-specific issues (19% overall; GitHub: 20%, Kaggle: 18%), data confusion (18% overall; GitHub: 17%, Kaggle: 19%), environment issues (16% overall; GitHub: 17%, Kaggle: 16%), and implementation errors (15% overall; GitHub: 15%, Kaggle: 15%).
-Finding 1.2: The main root causes of ML notebook crashes include API misuse (19%), notebook-specific issues (19%), data confusion (18%), environment problems (16%), and implementation errors (15%).
-Notebook-specific issues are a major cause of crashes, ranking first on GitHub and third on Kaggle.
-The majority of these crashes result from out-of-order execution (GitHub: 64%, Kaggle: 75%), where cells are executed in an incorrect sequence.
+3(b) visualizes the distribution of root causes for GitHub and Kaggle separately.
+The most common ones are API misuse (20.9% overall; GH: 21.0%, KG: 20.8%), notebook-specific issues (19.4%; GH: 19.2%, KG: 19.7%), implementation errors (16.9%; GH: 17.4%, KG: 16.3%), environment issues (16.8%; GH: 18.0%, KG: 15.5%), and data confusion (16.4%; GH: 16.9%, KG: 15.7%).
+Finding 1.3: The main root causes of ML notebook crashes include API misuse (20.9%), notebook-specific issues (19.4%), implementation errors (16.9%), environment problems (16.8%), and data confusion (16.4%).
+Notebook-specific issues are a major cause of crashes, ranking second on both GitHub and Kaggle.
+The majority of these crashes result from out-of-order execution (GH: 60.0%, KG: 74.3%), where cells are executed in an incorrect sequence.
 This occurs when users run notebook cells without considering dependencies.
-Another common cause is previous cell error (GitHub: 36%, Kaggle: 25%), where users encounter crashes in earlier cells but continue executing dependent cells without resolving the issues, resulting in additional crashes.
+Another common cause is previous cell error (GH: 40.0%, KG: 25.7%), where users encounter crashes in earlier cells but continue executing dependent cells without resolving the issues, resulting in additional crashes.
 The root cause distribution does not differ significantly between GitHub and Kaggle ($p>0.05$).
-Finding 1.3: Notebook-specific issues are the leading cause of crashes on GitHub ML notebooks and the third most common on Kaggle, with out-of-order execution being the most frequent (GitHub: 64%, Kaggle: 75%) problem.
-Root causes of exception types
-We merge both sample sets (GitHub and Kaggle) and plot the co-occurrence of exception types and their root causes in 4(a).
-Interestingly, we observe that root causes like API misuse, data confusion, and implementation errors contribute to a wide range of exceptions in ML notebooks.
-Notably, API misuse alone is responsible for at least 20% of crashes across multiple exception types, including ValueError (28%), TypeError (32%), APIArgumentError (63%), AttributeError (41%), KeyError (20%), and IndexError (27%).
-Meanwhile, data confusion contributes to over 30% of ValueError (43%), TypeError (35%), KeyError (55%), and IndexError (54%).
-These findings indicate that developers need to be attentive to the proper use of APIs and ensure that their data inputs are correctly formatted.
-Finding 1.4: API misuse and data confusion are the primary causes of a broad range of exception types.
-(a) Co-occurrence of exception types and root causes.
-(b) The percentage point (pp) difference, such that GitHub has $x$ pp more/fewer instances than Kaggle.
-Figure 4: Overview of the co-occurrence of crash exception types and root causes (see Table V for full names), and the difference in co-occurrence between Kaggle and GitHub.
-Certain exception types are dominantly linked to a specific root cause.
-For instance, IOError is connected to environment issues, APIArgumentError to API misuse, and KeyError and IndexError to data confusion.
-However, exceptions like ValueError, TypeError, and AttributeError have a wider range of root causes, indicating that these crashes may arise from more complex interactions in the code.
-Specifically, ValueError and TypeError can arise from data confusion (43%, 35%), API misuse (28%, 32%), and implementation errors (12%, 16%).
-AttributeError is primarily caused by API misuse (41%), implementation errors (20%), and data confusion (16%).
-Moreover, we find that 56% of NameError exceptions stem from notebook-specific issues which also account for 20% of KeyError occurrences.
-This highlights how the unique notebook semantics can introduce distinct challenges that manifest in particular exception types.
-Finding 1.5: ValueError, TypeError, and AttributeError arise from diverse root causes. NameError is mainly due to notebook-specific issues.
-4(b) highlights the absolute percentage differences in co-occurrence between GitHub and Kaggle notebook crashes.
-Barnard’s exact tests with Bonferroni correction show no significant difference in root cause distributions between GitHub and Kaggle for all exception types.
-However, some differences appear in the sample.
-Data confusion and API misuse both cause TypeError 1.6 times more commonly on Kaggle than on GitHub (a percentage points (pp) difference of 16).
-On GitHub, these are 4.5 and 1.8 times more commonly caused by intentional bugs and implementation errors (16 and 9 pp difference, respectively).
-Moreover, KeyError crashes are 1.5 times more commonly caused by data confusion on GitHub (23 pp difference), whereas notebook-specific issues are 3 times more frequent on Kaggle (18 pp difference).
-4.2 RQ2: To what extent are crashes in ML notebooks caused by ML bugs compared with general Python bugs?
+Finding 1.4: Notebook-specific issues are the second major cause of crashes in ML notebooks, with out-of-order execution being the most frequent (GH: 60.0%, KG: 74.3%).
+Root causes of crash types
+We find no statistical difference in root causes between GitHub and Kaggle, hence we merge them and plot crash type and root cause co-occurrences in Figure 4.
+We observe that API misuse, implementation errors, and data confusion are the root causes of a wide range of crashes in ML notebooks.
+Notably, API misuse alone is responsible for at least 25% of crashes observed along multiple crash types, including Invalid Argument (66%), Attribute Error (44%), Type Error (42%), Value Error (35%), Data Value Violation (31%), and Tensor Shape Mismatch (29%).
+Meanwhile, data confusion contributes to over 30% of several crash types, such as Key Error (50%), Data Value Violation (50%), Unsupported Broadcast (50%), Index Error (45%), Value Error (42%), Tensor Shape Mismatch (34%), and Feature Name Mismatch (33%).
+These findings indicate that debugging is likely to be difficult in such cases.
+Developers need to be attentive to the proper use of APIs and ensure that their data inputs are correctly formatted.
+Finding 1.5: API misuse, implementation errors, and data confusion are main causes of a wide range of crash types.
+Certain crash types are dominantly linked to one specific root cause.
+For instance, Out of Memory is always raised due to insufficient resources.
+IO Error, Environment Error, and Request Error are related to environment problems.
+Invalid Argument is mainly due to API misuse and data confusion.
+Name-related errors such as Variable Not Found, Name Error, and Module Not Found are mostly connected to notebook-specific issues.
+This highlights how notebook semantics can introduce unique challenges that manifest in particular crash types.
+Finding 1.6: Name-related errors, in particular Variable Not Found, are mainly due to notebook-specific issues.
+To assess the accuracy of our root cause analysis, we try to reproduce 30 crashes in Kaggle and successfully reproduce 20.
+We then resolve the crashes, finding that 18 ($90\%$) are caused by the root causes identified in this study.
+This high accuracy suggests that our approach is reliable.
+Figure 4: Overview of the co-occurrence of crash types and root causes (see Table V for full names).
+4.2 RQ2: To what extent are crashes in ML notebooks caused by ML bugs and by which ML libraries?
 Crashes in GitHub and Kaggle notebooks can stem from either code that uses ML libraries or general Python code.
@@
 Additionally, we examine the root causes of both ML and Python bugs, as shown in Figure 5, to gain deeper insights into their distinctions and similarities.
-ML bugs and Python bugs
-ML bugs account for 64% of crashes overall in ML notebooks, with 59% on GitHub and 71% on Kaggle.
-The $\chi^{2}$ test reveals a significant difference in the distribution of ML and Python bugs between the two platforms ($p<0.001$).
-These findings suggest that a substantial proportion of crashes on both platforms stems from code that directly interacts with ML libraries, with the impact being particularly pronounced in Kaggle notebooks.
-Finding 2.1: A higher percentage of crashes in both GitHub and Kaggle notebooks are attributed to ML bugs (GitHub: 59%, Kaggle: 71%) compared to Python bugs.
-Kaggle exhibits an even more significant prevalence of ML-related crashes than GitHub.
+Furthermore, we examine the frequency of crashes directly linked to the use of ML libraries (see Figure 6) to identify which libraries present the greatest challenges.
+ML bugs vs. Python bugs
+ML bugs account for 67.7% of crashes overall in ML notebooks (GH: 62.8%, KG: 73.0%).
+The $\chi^{2}$ test reveals a significant difference in the distribution of ML and Python bugs between the two platforms ($p<0.01$).
+These findings suggest that a substantial proportion of crashes on both platforms stems from code that interacts with ML libraries, with the impact being particularly pronounced in Kaggle notebooks.
+Finding 2.1: On both platforms, we find significantly more ML bugs than Python bugs. Kaggle exhibits a more significant prevalence of ML-related crashes than GitHub.
 Root causes of ML and Python bugs
-As Figure 5 illustrates, for ML-related crashes, the primary causes are API misuse (23%), data confusion (23%), and notebook-specific issues (21%).
-Environment setup problems (12%) and implementation errors (10%) are also big factors.
-In contrast, general Python crashes are mainly driven by environment issues (24%) and implementation errors (24%), followed by notebook-specific issues (17%) and API misuse (13%).
-Finding 2.2: The primary causes of ML-related crashes are API misuse (23%), data confusion (23%), and notebook-specific issues (21%), while general Python crashes are mostly caused by environment problems (24%), implementation errors (24%), and notebook-specific issues (17%).
+As Figure 5 illustrates, the primary causes of ML-related crashes are API misuse (25%), notebook-specific issues (20%), and data confusion (20%).
+Environment setup problems (12%) and implementation errors (11%) are also big factors.
+Furthermore, ML-specific bugs include unique root causes such as ML model mismatch (3%), insufficient resources (3%), and library issues (1%).
+In contrast, general Python crashes are mainly driven by implementation errors (29%) and environment issues (28%), followed by notebook-specific issues (18%) and API misuse (13%).
+The differences in root causes between ML and general Python bugs stem from the inherent complexity of working with ML models and data, as well as the common use of notebook semantics when performing ML tasks.
+While the saved states in notebooks improve usability, they can also increase the likelihood of crashes if not properly managed.
+Finding 2.2: The primary causes of ML-related crashes are API misuse (25%), notebook-specific issues (20%), and data confusion (20%), while general Python crashes are mostly caused by implementation errors (29%) and environment problems (28%).
 Figure 5: Overview of the overall co-occurrence of ML/Python bugs and root causes (see Table V for full names).
 The root cause distributions between ML and Python bugs differ significantly ($p<0.001$).
-Notably, the higher percentage of API misuse in ML-related crashes (23%) compared to general Python crashes (13%) underscores the challenges developers face in correctly using ML libraries.
+Notably, the higher percentage of API misuse in ML-related crashes (25%) compared to general Python crashes (12%) underscores the challenges developers face in correctly using ML libraries.
 We find no statistically significant differences between the platforms in the distribution of root causes for either ML bugs or Python bugs ($p>0.05$ for both).
-Finding 2.3: API misuse is more prevalent in ML-related crashes (23%) than in general Python crashes (13%).
-4.3 RQ3: In which stages of the ML pipeline are crashes in notebooks most likely to occur?
-The ML pipeline outlines key stages of ML programs.
-We analyze the frequency of crashes occurring in these stages during ML development in notebooks and determine how often crashes in a certain stage involve ML bugs versus general Python bugs.
+Finding 2.3: API misuse is more prevalent in ML-related crashes (25%) than in general Python crashes (12%).
+Figure 6: The challenge factors ($CF$) of the top 10 ML libraries, indicating the ratio between a library’s involvement in crashes and its usage. See Table II for library descriptions.
+ML libraries
+Figure 6 ranks the top 10 most commonly used ML libraries by their challenge factor ($CF$), which measures how likely using a ML library is to cause crashes compared to how often it is used.
+The Challenge factor is computed as $CF=Lib_{\text{crashes}}/Lib_{\text{usage}}$, where 1) $Lib_{\text{crashes}}=N_{\text{Lib\_crash}}/N_{\text{crashes}}$ is the proportion of crashes caused by using an ML library.
+2) $Lib_{\text{usage}}=N_{\text{Lib\_use}}/N_{\text{notebooks}}$ represents how frequently the library appears in ML notebooks, counting each notebook only once per library.
+A higher $CF$ indicates a library is more crash-prone relative to its usage.
+In this ranking, we merge tensorflow and keras because keras has been tightly integrated into tensorflow since TensorFlow 2.0.
+Combining them provides a more accurate representation of their impact compared to other ML libraries.
+ML libraries tensorflow/keras, and torch exhibit the highest challenge factors (over 5), followed by sklearn (2.8), pandas (2.5), and torchvision (1.3).
+Finding 2.4: tensorflow/keras and torch are the top ML libraries most challenging to use correctly.
+4.3 RQ3: In which stages of the ML pipeline are crashes in notebooks most frequently to occur?
+The ML pipeline outlines the key stages of ML programs.
+We analyze the frequency of crashes occurring in these stages during ML development in notebooks and determine how often crashes in a certain stage involve ML bugs.
 The results are shown in Figure 7.
 ML pipeline
-6(a) shows that most crashes occur during the data preparation (27% overall; GitHub: 29%, Kaggle: 26%) stage, model training (16% overall; GitHub: 13%, Kaggle: 19%), evaluation/prediction (15% overall; GitHub: 11%, Kaggle: 21%), and data visualization (11% overall; GitHub: 10%, Kaggle: 12%) stages.
-Finding 3.1: The majority of crashes occurs during data preparation (27%), model training (16%), and evaluation/prediction (15%) stages in the ML pipeline.
-Notably, 25% of crashes on GitHub are unrelated to any pipeline stage, which is approximately 3.5 times more than Kaggle (7%).
-We apply the $\chi^{2}$ test to evaluate this, revealing a significant difference in distributions ($p<0.001$).
+7(a) shows that most crashes occur during the data preparation (33.0% overall; GH: 37.7%, KG: 27.8%) stage, model training (19.4%; GH: 17.4%, KG: 21.6%), evaluation/prediction (18.4%; GH: 14.6%, KG: 22.5%), and data visualization (13.5%; GH: 14.1%, KG: 12.9%) stages.
+Crashes during data preparation are likely due to the complexities of cleaning and processing datasets.
+Model training and evaluation crashes may result from incorrect settings, misaligned inputs, or resource limitations.
+Data visualization errors are likely caused by mismatched data types or improper library usage.
+Finding 3.1: The majority of crashes occurs during data preparation (33.0%), model training (19.4%), and evaluation/prediction (18.4%) stages in the ML pipeline.
 (a) ML pipeline stages
@@
 Figure 7: Overview of crashes across the ML pipeline (see Table V for full names) with the distribution of ML/Python bugs per stage and the adjusted residuals per stage.
-Finding 3.2: GitHub notebooks show 3.5 times more crashes than Kaggle not related to ML pipeline stages.
-We then examine differences between GitHub and Kaggle in ML pipeline stages, excluding crashes unrelated to any stage.
+We then examine the differences between GitHub and Kaggle in the ML pipeline stages.
 A $\chi^{2}$ test reveals a significant difference ($p<0.01$).
-To understand which stages contribute more to this difference, we calculate the adjusted residuals [34], comparing observed crashes in each stage with the expected number of crashes under the null hypothesis (shown in 6(b)).
+To understand which stages contribute more to this difference, we calculate the adjusted residuals [39], comparing observed crashes in each stage with the expected number of crashes under the null hypothesis (shown in 7(b)).
 A larger absolute residual indicates a greater contribution.
-The results suggest that GitHub contains more crashes in the environment setup and data preparation, and Kaggle in the evaluation/prediction stage.
-Finding 3.3: GitHub and Kaggle crashes are distributed differently across ML pipeline stages, especially in environment setup, data preparation, and evaluation/prediction.
+The results suggest that GitHub contains more crashes in the environment setup and data preparation, likely due to the variability in user environments and dependencies.
+Kaggle, with pre-configured environment, faces more crashes during evaluation/prediction, possibly due to model or input setting issues, or resource limitations.
+Finding 3.2: GitHub and Kaggle crashes are distributed differently across ML pipeline stages.
+GitHub has more during environment setup and data preparation, whereas Kaggle sees more during the evaluation/prediction stage.
 ML/Python bugs per ML pipeline stage
-The breakdown of ML and Python bugs in each pipeline stage reveals that ML bugs dominate critical stages, with over 80% of crashes during model training and construction and over 70% during evaluation/prediction for both platforms.
-Finding 3.4: ML-related crashes are dominant (over 70%) in model training, construction, and evaluation/prediction stages of the ML pipeline.
-We apply Barnard’s exact test to assess significant differences of ML/Python bug distributions between GitHub and Kaggle for each ML pipeline stages.
-The results show significant differences in the environment setup ($p<0.05$) and data preparation ($p<0.01$) stages.
+The breakdown of ML and Python bugs in each pipeline stage reveals that ML bugs dominate critical stages, with over 80% of crashes during model construction and training, and over 70% during evaluation/prediction for both platforms.
+Finding 3.3: ML-related crashes are dominant (over 70%) in model construction, training, and evaluation/prediction stages of the ML pipeline.
+We apply Barnard’s exact test to assess significant differences of ML/Python bug distributions between GitHub and Kaggle for each ML pipeline stage.
+The results show significant differences in the data preparation ($p<0.01$) stage.
 In Kaggle, 70% of crashes during data preparation are ML-related, whereas in GitHub, ML and Python bugs occur at similar rates.
-In the environment setup stage, most Kaggle crashes are due to Python bugs, only 16% are ML bugs, while in GitHub, 45% of crashes involve ML bugs.
-Finding 3.5: Kaggle has 1.4 times more ML-related crashes in the data preparation stage than GitHub, whereas GitHub shows almost 3 times more in the environment setup stage.
+Finding 3.4: Kaggle has 1.4 times more ML-related crashes in the data preparation stage than GitHub.
 Root causes of ML pipeline stages
 We further analyze root cause distribution across ML pipeline stages (see Figure 8).
-Most stages display a range of root causes, including API misuse, notebook-specific issues, data confusion, environment issues, and implementation errors.
+Most stages display a range of root causes, including API misuse, notebook-specific issues, implementation errors, environment issues, and data confusion.
 Notably, in the environment setup stage, 84% of crashes stem specifically from environment-related issues.
@@
 Notebook-specific issues
-As discussed in RQ1 (Section 4.1), notebook-specific issues are a major cause (19%) of crashes, contributing to 56% of NameError and 20% of KeyError exceptions.
-This highlights how notebook semantics can become a liability, increasing the risk of crashes.
-For instance, if a variable x is defined in one cell but a later cell that references x is executed first, a NameError occurs due to out-of-order execution since x has not been initialized yet.
-Similarly, if a developer incorrectly tries to add a column to a dataframe, causing the cell to crash, and then runs another cell that references this non-existent column without first correcting the initial error, a KeyError will occur due to previous cell error.
-The flexibility to execute cells in any order can easily disrupt cell dependencies, especially in complex ML notebooks.
-This underscores the unique challenges of bug detection and debugging in notebooks, which requires careful tracking of variable states and cell execution order at the current execution point.
-NBSAFETY [35] addresses this issue by providing a custom Jupyter kernel that flags potentially unsafe cell executions.
-Takeaway 1: The flexibility of the notebook environment can cause instability, as out-of-order execution often leads to crashes like NameError.
+As discussed in RQ1 (Section 4.1), notebook-specific issues contribute to a significant portion (19.4%) of crashes, particularly those related to name resolution (e.g., Variable Not Found, Name Error, Module Not Found), Model Initialization Error, and key errors.
+These results highlight how the flexible execution model of notebooks can introduce instability, increasing the likelihood of crashes.
+Unlike conventional software engineering workflows, where execution order is typically linear and well-defined, notebooks enable an interactive, fragmented development process.
+This flexibility, while beneficial for experimentation, also disrupts state management, leading to errors caused by out-of-order execution and implicit dependencies.
+For instance, if a variable x is defined in one cell but a later cell that references x is executed first, a Variable Not Found occurs due to out-of-order execution since x has not been initialized yet.
+These issues underscore the unique challenges of debugging in notebooks, where errors may not be immediately apparent and can propagate unpredictably.
+This finding underscores notebook development challenges, particularly in managing hidden state and execution order dependencies.
+Tools like NBSAFETY [40] attempt to mitigate these risks by providing a custom Jupyter kernel that detects unsafe execution sequences.
+However, our results suggest that the problem remains prevalent, reinforcing the need for improved debugging support tailored to the dynamic execution model of notebooks.
+Takeaway 1: The flexibility of notebooks disrupts state management, leading to unusual execution-order errors, which highlight the need for better debugging tools to support ML developers.
 Data confusion
-Another common root cause of crashes is data confusion (18%), indicating practical challenges in handling heterogeneous data types.
-It is a key contributor to a variety of exceptions such as ValueError, TypeError, KeyError, and IndexError, indicating that tackling data confusion could help resolve many of these exceptions.
-Data confusion also accounts for a high percentage of the ML-related crashes (23%), as discussed in RQ2 (Section 4.2).
-It suggests that improper data handling such as incorrect data manipulations for ML APIs, is critical in ML-specific code.
-Specifically, ML development often requires transformations of data into specific structures that can be easily mishandled, causing errors like shape mismatches, improper data transformations, or incorrect assumptions about data formats.
-This underscores the need for improved data validation.
-Takeaway 2: Data handling poses a big challenge when using ML notebooks.
+Another common root cause of crashes is data confusion (16.4%), indicating practical challenges in handling heterogeneous data types.
+There are many crash types, such as Key Error, Data Value Violation, Unsupported Broadcast, and Index Error, which point to the same root cause of data confusion.
+This potentially complicates fault localization process.
+Data confusion also accounts for a high percentage of ML-related crashes (20%) (see Section 4.2), emphasizing the impact of improper data handling, such as incorrect transformations for ML APIs.
+Specifically, ML development often requires precise data formatting, and errors like shape mismatches or faulty transformations can be hard to allocate.
+This underscores the need for better data validation tools to prevent such issues.
+Takeaway 2: Data confusion in ML notebooks leads to diverse crash types, making fault localization challenging.
 ML bugs
-When analyzing the results for RQ2 (Section 4.2), we find that crashes in ML notebooks are overall 1.8 times more frequently linked to ML bugs compared to Python bugs, reflecting key differences between developing ML and traditional software.
-Notably, a significantly higher proportion of ML-related crashes (23%) are directly attributed to API misuse, compared to only 13% for general Python crashes.
-The extensive reliance on ML libraries increases API misuse errors, particularly when developers are unfamiliar with libraries’ specific requirements.
-Although these libraries are crucial for effective data processing and modeling, they often have complex usage guidelines [36, 37].
-Misinterpretations of parameter settings, or input-output expectations frequently result in crashes.
-To mitigate these issues, methods that verify API usage and enhance error-handling mechanisms could be beneficial.
-Takeaway 3: Correct use of ML libraries is challenging in ML notebook development, highlighting the importance of verifying API usages and enhancing error-handling.
+When analyzing the results for RQ2 (Section 4.2), we find that crashes in ML notebooks are overall 2.1 times more frequently linked to ML bugs compared to Python bugs, reflecting key differences between developing ML and traditional software.
+Notably, a significantly higher proportion of ML-related crashes (25%) are attributed to API misuse, compared to only 13% for general Python crashes.
+This issue stems from the heavy reliance on ML libraries, with complex usage guidelines [41, 42] that often lead to misinterpretations of parameters and input-output expectations, causing crashes.
+Our analysis reveals that certain DL libraries, particularly tensorflow/keras and torch, pose greater challenges.
+In tensorflow/keras, tensor shape mismatches are a major cause of crashes, often due to implicit broadcasting, dynamic input shapes, or insufficient safeguards for detecting misaligned dimensions before execution.
+In contrast, torch is more prone to runtime errors, such as device compatibility issues (e.g., loading CUDA models on a CPU) and incorrect loss function configurations.
+Some of these errors may arise from unclear documentation, while others may stem from the flexibility of these libraries, where errors only surface at runtime, making static checks difficult.
+More informative error messages and warnings could significantly aid debugging.
+These findings suggest that many crashes may stem not just from user mistakes but also from library design choices, including complex APIs, ambiguous error messages, and limited safeguards.
+This highlights the need for library developers to improve API usability, enhance documentation clarity, and implement better error-handling mechanisms.
+Additionally, identifying error-prone libraries can help practitioners prioritize debugging efforts, develop better tools, and choose libraries that align with their stability needs.
+Takeaway 3: ML notebooks are prone to crashes due to API misuse, especially when using deep learning libraries tensorflow/keras and torch.
 The data preparation stage
-RQ3 (Section 4.3) demonstrates that most ML notebook crashes occur during the data preparation stage (GitHub: 29%, Kaggle: 26%) of the ML pipeline.
+RQ3 (Section 4.3) demonstrates that ML notebook crashes occur mostly during the data preparation stage (33.0%) of the ML pipeline.
 This highlights the primary challenge developers face: preparing or preprocessing datasets to ensure compatibility with model requirements.
 Issues often arise from understanding the data itself or aligning it with API expectations.
-Additionally, the limited support available for developers during this stage implies they may spend a disproportionate amount of time here.
+Additionally, the limited support available for developers during this stage implies that they may spend a disproportionate amount of time here.
 These insights suggest that improving tooling and resources for data preparation could significantly enhance developer productivity.
@@
 The model training and evaluation/prediction stages
-Crashes are more common in the model training (16%) and evaluation/prediction (15%) stages than in the model construction (7%) stage.
-These errors typically arise at runtime, when models are trained or evaluated with actual data inputs, rather than during the construction phase when input shapes are often flexible.
-Libraries like TensorFlow, for instance, may not fully validate data compatibility until training begins, which can lead to errors surfacing only in these later stages.
+Crashes are more common in the model training (19.4%) and evaluation/prediction (18.4%) stages than in the model construction (8.3%) stage.
+These errors typically arise at runtime, when models are trained or evaluated with actual data inputs, rather than during the construction phase, when input shapes are often flexible.
+Libraries like tensorflow/keras, for instance, may not fully validate data compatibility until training begins, which can lead to errors surfacing only in these later stages.
 When issues occur within specific model layers, error messages can be ambiguous, making it difficult for developers to pinpoint and resolve problems.
 Consequently, developers may need to iteratively debug by re-training and re-evaluating multiple times, a tedious and time-intensive process.
-Notebook semantics help preserve previous cell executions, so developers can avoid re-running all preceding cells (e.g., data preparation).
-However, this approach also increases the risk of crashes from, for example, out-of-order execution, where dependencies may no longer align.
+Developers can avoid re-running all preceding cells as notebook semantics preserve previous cell executions, speeding up the iterative debugging process.
+However, this also increases the risk of crashes from, for example, out-of-order execution, where dependencies may no longer align.
 These challenges highlight the importance of enabling early bug detection without requiring code execution for training or evaluating ML models, along with mechanisms to track cell dependencies and maintain the correct execution order.
@@
 Comparing ML notebook crashes between GitHub and Kaggle reveals their key differences.
-GitHub has a significantly higher proportion of crashes outside typical ML pipeline stage (25% vs.
-7% on Kaggle), suggesting a more diverse use of notebooks where crashes often occur outside standard ML workflows.
+In the cases excluded from the initial sample, GitHub has a significantly higher proportion of crashes outside typical ML pipeline stages (25% vs.
+7% on Kaggle), for example, tutorial notebooks, educational materials, or physics simulations.
+This suggests a more diverse use of notebooks where crashes often occur outside standard ML workflows.
 Moreover, there is a notable difference in crash distributions across ML pipeline stages between the two platforms.
-For example, GitHub has more crashes during environment setup, likely due to its user-driven environments that can lead to inconsistencies in software versions or dependencies.
-GitHub also sees more crashes in data preparation, likely because it often uses “raw” datasets that require more complex preprocessing, whereas Kaggle provides standardized datasets with detailed descriptions.
+For example, GitHub sees more crashes in data preparation, likely because it often uses “raw” datasets that require more complex preprocessing, whereas Kaggle provides standardized datasets with detailed descriptions.
 In contrast, Kaggle has more crashes during evaluation/prediction, possibly due to its submission process for competitions, which introduces additional model evaluation complexities.
 ML/Python bug
-Kaggle shows a significantly higher prevalence of ML-related crashes than GitHub (71% vs.
-59%).
-One exception is during the environment setup stage of the ML pipeline, where GitHub shows 3 times more ML-related crashes.
+Kaggle shows a significantly higher prevalence of ML-related crashes than GitHub (73.0% vs.
+62.8%).
+One exception is during the environment setup stage of the ML pipeline, where GitHub shows 2.6 times more ML-related crashes.
 This is likely because Kaggle’s pre-configured environment includes essential ML libraries, reducing setup-related errors.
@@
 5.3 Comparison with existing work
-Frequent bugs in notebooks
+To contextualize our results, we compare our findings with previous studies [13, 10].
+Zhang et al.
+[13] identify failure types for DL jobs similar to our crash types.
+They also report the distribution of failures across ML pipeline stages.
 De Santana et al.
-[8] find that environment and configuration issues are the most common bugs in Jupyter notebooks, often discussed on platforms like Stack Overflow.
-However, our study shows a higher prevalence of coding errors (52%), including API misuse, data confusion, and implementation issues.
-This difference likely reflects data sources: while the analysis from De Santana et al.
-focuses on user-reported issues, our study examines error outputs directly from executed notebooks.
-Our approach captures runtime failures, highlighting coding-related errors, whereas environment issues are often resolved before execution and thus less visible in our dataset.
-ML/Python bugs
-Our study finds that ML crashes are more frequent than general Python crashes (Kaggle: 71%, GitHub: 59%), in contrast with findings by Zhang et al.
-[11].
-This discrepancy arises partly from differing definitions: they define ML bugs at a system level, encompassing all bugs in ML components of a system, while we focus specifically on bugs in code interacting with ML libraries.
-Additionally, both studies analyze error messages, but theirs are derived from job failures where code is tested, whereas ours come from cell outputs in notebooks, capturing issues during early development.
-This suggests that our study more directly reflects initial challenges faced by developers.
-ML pipeline
-Our study finds that the data preparation stage is the most crash-prone (GitHub: 29%, Kaggle: 26%) in the ML pipeline, similar to Islam et al.
-[7].
-While their study includes bugs related to poor performance and incorrect outputs, not just crashes, and does not detail the impact of bugs within each stage, both studies highlight that data preparation remains the most challenging stage.
-Root cause identification
-Our study identifies new root causes of crashes in ML notebooks, particularly those related to notebook-specific issues like out-of-order execution and previous cell error.
-While out-of-order execution has been linked to reproducibility issues in prior work [2], it has not been explored as a crash root cause.
-Additionally, by analyzing crash information, context code, and cell outputs, we can distinguish crashes caused by data confusion from API misuse.
-Error messages and cell outputs reveal details about data shapes, types, and structures that reflect users’ understanding of their data, helping us identify crash origins.
-Examining cell outputs and documentation cells also enables us to identify intentional errors, such as API misuse shown in tutorial notebooks for instructional purposes.
+[10] study bugs in Jupyter notebooks at a general level, which are similar to our target data format.
+To enable a fair comparison, we map our results to existing classifications of failure types [13] and bug types [10].
+This is unnecessary for ML pipeline stages as our categorization includes the categories used in the other study [13].
+For the study by Zhang et al.
+[13], we retain only DL notebooks to match their domain.
+This filtering process follows a similar approach to identifying ML notebooks, as described in Section 3.2, where we compile a list of DL libraries as shown in Table II.
+Following the same methodology as ML library identification, two authors independently label the DL libraries, achieving a 100% agreement rate.
+The overall mapped results are presented in Table VII, while details on our mapping strategy and subcategory comparisons are available in our GitHub repository [15].
+Bug type comparison with Zhang et al. [13]
+As VII(a) shows, execution environment errors are substantially lower in our DL dataset (compared to Zhang et al.
+[13]).
+The authors attribute the high frequency of execution environment errors to differences between job execution environments and engineers’ local environments.
+Our findings indirectly support this claim.
+Our target artifacts are DL notebooks typically executed either locally (GitHub) or in controlled environments with pre-installed dependencies (Kaggle).
+When developing DL notebooks in such stable environments, environment-related bugs are less common.
+Transitioning to other environments significantly increases the likelihood of encountering environment errors.
+Takeaway 7: Crashes caused by execution environment are less frequent in ML notebooks but they can emerge when transitioning to execution platforms (e.g., cloud-based) used for deep learning.
+Practitioners should proactively test for such issues to ensure smooth deployment.
+After excluding the effect of execution environment errors, our study still identifies a higher prevalence of DL-specific crashes.
+This discrepancy can be attributed to the different development stages being analyzed.
+Our results derive from examining cell outputs in notebooks primarily used for prototyping and early-stage DL development.
+In contrast, Zhang et al.
+mainly focus on later development stages when jobs (tested locally) are already being submitted to a DL platform.
+As such, our study should better represent the initial difficulties faced by DL developers.
+Takeaway 8: As ML notebooks are used in early stage of development, deep learning-specific crashes are more prevalent in ML notebooks compared to deep learning jobs.
+Bug type comparison with De Santana et al. [10]
+As shown in VII(b), De Santana et al.
+[10] find that environment and settings errors are considerably more common in Jupyter notebooks, as observed in GitHub commits, compared to our findings.
+This discrepancy is likely due to differences in the nature of the analyzed notebooks.
+Specifically, their dataset primarily consists of notebooks from stable and mature GitHub projects, such as AWS, Azure, Fastai, and TensorFlow, whereas our dataset focuses on notebooks used for ML prototyping and experiments.
+To better understand this difference, we randomly sample 20 commits from their dataset to observe that the associated notebooks are predominantly demonstrations, documentation, or usage examples of specific libraries or software created by the developers of those libraries.
+The corresponding commits are mostly aimed at updating dependencies required by the projects rather than fixing implementation bugs, reflecting the stable and mature nature of these projects.
+As a result, it is expected that environment and settings issues are more prevalent in their dataset.
+We believe that our study better reflects the practical development challenges in notebook environments.
+We observe unique implementation errors, particularly those related to notebook-specific issues like out-of-order execution and previous cell errors.
+While out-of-order execution has been linked to reproducibility concerns in prior research [2], its role as a root cause of crashes has not been widely explored.
+By analyzing crash information, code, and cell outputs, our study also distinguishes crashes caused by data confusion from those resulting from API misuse.
+Error messages and outputs provide crucial insights into data shapes and structures, helping us pinpoint the origin of crashes more effectively.
+Therefore, we present these implementation issues separately in VII(b) by marking them as “Implementation (other).”
+(a) Bug type classification comparison with Zhang et al.
+(S1 [13]).
+The last column shows normalized percentages excluding execution environment (Env.).
+Our results are based on DL-only notebooks crashes.
+Dimensions S1 [13] Result S1 [13] Result Ours (DL) Exclude Env. S1 [13] vs. Ours
+Execution environment 48.0% 19.9% -
+Data 2.0% 0.9% 3.8% vs. 1.1%
+DL specific 13.5% 32.5% 26.0% vs. 40.6%
+General code error 36.5% 46.7% 70.2% vs. 58.3%
+(b) Bug type classification comparison with De Santana et al. (S2 [10]). We present our mapped results separately for Kaggle and GitHub for fair comparison.
+Dimensions S2 [10] Result S2 [10] (GH) Result Ours (KG) Result Ours (GH)
+Environment and settings 35.6% 3.4% 6.9%
+Processing 1.9% 3.1% 1.5%
+Implementation 44.2% 26.4% 23.1%
+Implementation (other) - 67.1% 68.5%
+Other (e.g., Kernel) 18.3% - -
+(c) ML pipeline stage comparison with S1 [13] (DL cases only). Our normalized percentages exclude data visualization (DV) and model construction (MC) stages absent in S1.
+ML pipeline stages Our category Result S1 [13] Exclude DV&MC Ours (DL) Result Ours (DL)
+Environment setup 30.8% 10.0% 7.6%
+Data preparation 30.0% 22.7% 17.3%
+Data visualization - - 7.6%
+Model construction - - 16.4%
+Training 15.0% 35.7% 27.2%
+Evaluation/prediction 24.3% 31.6% 24.0%
+TABLE VII: Comparison with existing studies [13, 10]. An overview of these studies is provided in Table I.
+Takeaway 9: The error patterns prevalent in Jupyter notebooks created by authors of ML libraries are different from ML notebooks created by the users of such libraries.
+ML pipeline comparison with Zhang et al. [13]
+As VII(c) shows, Zhang et al.
+[13] report considerably more environment setup failures but fewer training failures than our results.
+Their higher environment setup failure rate aligns with their larger proportion of execution environment bugs (VII(a)).
+Fewer training errors may reflect that DL developers focus more on ensuring training correctness when developing locally before job submission.
+As a novel finding, our analysis reveals that data visualization can also be a source of crashes in ML workflows because notebooks serve not only for coding but also for documentation and exploratory data analysis.
+The omission of this stage in prior work suggests that visualization-related crashes might be underrepresented.
+Identifying this category highlights the unique nature of notebook development and underscores the need for debugging support tailored to interactive workflows.
+Takeaway 10: Since ML notebooks blend coding with exploration and documentation, notebook development has some unique types of errors uncommon in traditional ML workflows, particularly, in data visualization.
 6 Threats to validity
@@
 6.1 External validity
-Temporal limitation in GitHub dataset
-The GitHub dataset includes notebooks from January 1, 2015, to March 31, 2022, which may not fully represent current notebook trends.
-Given the rapid evolution of coding practices and the tools used over time, notebooks from earlier years may reflect outdated coding styles or libraries.
-To mitigate this, we apply a filtering process to remove notebooks using older Python versions, improving the generalizability of the dataset.
-Comparison of GitHub and Kaggle notebooks
-The Kaggle dataset, which spans from January 1, 2023, to January 1, 2024, provides a more up-to-date snapshot of notebook trends.
-While the differing periods of GitHub and Kaggle datasets reduce redundancy in notebooks and associated crashes, the validity of comparison is threatened because ML practices may have changed.
-However, there is limited evidence of this effect because our results show no significant difference in exception types and root causes between Kaggle and GitHub.
+Data selection bias
+As our dataset only includes publicly shared GitHub and Kaggle notebooks, it may exclude crashes that users fix before publishing, thus crashes encountered during development may be underrepresented.
+As a result, our findings reflect crashes that persist, which may not be the full spectrum of debugging challenges.
+Future work could address this by analyzing version histories or autosaved checkpoints to capture a more complete development process.
+Temporal limitations in datasets
+The GitHub dataset includes notebooks from January 1, 2015, to March 31, 2022, while the Kaggle dataset spans from January 1, 2023, to January 1, 2024.
+Given the rapid evolution of coding practices and tools used over time, notebooks from earlier years may reflect outdated coding styles or libraries.
+To mitigate this, our filtering process removes notebooks using older Python versions, improving the generalizability of the dataset.
 KeyboardInterrupt exceptions
@@
 6.2 Internal validity
-Tutorial notebooks
-The analyzed dataset contains various tutorial and educational notebooks.
-These tutorial notebooks often contain intentional crashes for didactic purposes.
-Such crashes could distort the crash patterns we aim to uncover.
-However, while these crashes are intentional, they still represent errors that users commonly believe occur in real-world scenarios, thus minimally affecting our results.
+Clustering crashes
+We cluster crashes with similar patterns based on their error values.
+We enhance this by removing specific substrings, like key names and numeric values.
+However, this results in an empty string in cases where error values only contain these specific substrings, which are consequently grouped into the same cluster.
+Examples are KeyError and SystemExit, which are very different errors.
+Although a theoretical threat, an investigation indicates that it only accounts for 0.38% of the dataset.
+Sampling limitations
+Our proportional stratified sampling ensures broad representation, but rare strata may still be underrepresented despite a minimum sample size of one per stratum.
+Increasing this minimum was impractical due to manual labeling constraints.
+To assess potential gaps, we conduct an exploratory analysis by manually examining a maximum of 10 crashes from 10 randomly selected small clusters in GitHub and Kaggle datasets separately.
+This reveals no previously unknown crash patterns, suggesting that our approach captures key crash types and root causes, although some rare errors may still be missed.
+Furthermore, cases excluded during resampling reduce certain strata sizes, which may impact the original sampling proportions.
 Manual labeling
@@
 Any disagreements are resolved through consensus.
-Reproducibility of crashes
-The error outputs we analyze represent real execution crashes but they are not necessarily reproducible due to factors like execution order, missing datasets, or environment replication [2].
+Reproducibility of crashes and root cause analysis
+The error outputs we analyze represent real execution crashes, but they are not necessarily reproducible due to factors like execution order, missing datasets, or environment replication [2].
 While reproducing crashes could further validate our findings, it is challenging, especially for GitHub notebooks.
 Kaggle notebooks, with their accessible datasets and pre-configured environments, may enable reproducibility.
-In future work, we plan to reproduce crashes in Kaggle notebooks to validate our results and create a reproducible dataset for deeper analysis.
-Clustering crashes
-We cluster crashes with similar patterns based on their error values.
-We enhance this by removing specific substrings, like key names and numeric values.
-However, this results in an empty string in cases where error values only contain these specific substrings, which are consequently grouped into the same cluster.
-Examples are KeyError and SystemExit, which are very different errors.
-Although a theoretical threat, an investigation indicates that it only accounts for 0.38% of the dataset.
+Nevertheless, we carried out a small-scale validation to assess the accuracy of our root cause analysis by trying to reproduce a few crashes in Kaggle, and the high accuracy suggests that our approach is reliable (see Section 4.1).
+In future work, we plan to extend crash reproducibility efforts in Kaggle notebooks to validate our results thoroughly and create a reproducible dataset for deeper analysis.
 7 Conclusions and Future Work
-In this paper, we present the first focused study on software crashes in ML notebooks from GitHub and Kaggle, uniquely identifying notebook-specific root causes, such as out-of-order execution and previous cell error.
-Our analysis classifies exception types, root causes, and ML pipeline stages, distinguishing between ML-specific and general Python bugs.
-Our findings reveal that the most common exceptions in ML notebooks are NameError, ValueError, and TypeError.
-The primary root causes are API misuse, data confusion, and notebook-specific issues (especially out-of-order execution).
-Notably, ML bugs account for more crashes than Python bugs, especially in Kaggle notebooks.
-Most crashes (58%) occur during the data preparation, model training, and evaluation/prediction stages of the ML pipeline.
-Additionally, GitHub and Kaggle exhibit different crash distributions across ML pipeline stages.
-Future work will expand on these findings by constructing a reproducible dataset of ML crashes.
-Motivated by the results of this study, we also aim to develop an integrated, real-time bug detection tool tailored to ML development in the notebook environment, building on our previous work [39].
-This tool will focus on ML bug detection while monitoring cell execution order and maintaining notebook semantics, helping developers catch errors as they code and enhancing productivity in ML development in notebooks.
+This paper provides the first large-scale empirical study focusing on software crashes in public ML notebooks collected from GitHub and Kaggle.
+Our analysis classifies (1) crash types, (2) root causes, (3) ML pipeline stages, and (4) involved ML libraries.
+(5) We also distinguish between ML-specific and general Python bugs.
+Our findings reveal that the (1) most common crashes in ML notebooks are Variable Not Found, Invalid Argument, and IO Error.
+We have also identified prevalent ML-specific crash types such as Tensor Shape Mismatch, Data Value Violation, and Out of Memory.
+Moreover, (2) the primary root causes of crashes are API misuse and notebook-specific issues (specifically, out-of-order execution and previous cell error).
+(3) Most crashes (70.8%) occur during data preparation, model training, and evaluation/prediction stages of the ML pipeline, while data visualization bugs are observed for the first time.
+(4) We also identify that certain ML libraries are more challenging to use for ML notebook development, in fact, crashes mostly occur when using DL libraries tensorflow/keras and torch.
+Finally, (5) ML bugs account for more crashes than Python bugs, especially in Kaggle notebooks.
+Future work will expand on these findings by constructing a reproducible dataset of ML crashes, and investigate the temporal evolution of crash patterns.
 Acknowledgments

### v3

@@
 In this context, Jupyter notebooks, particularly Python-based ones, have been widely adopted for ML prototyping and data analysis [6].
-In industrial settings [7], Jupyter notebooks are often the starting point for developing ML applications before transitioning to production-ready ML pipelines, where much of the Python code used in notebooks is reused in large-scale distributed computing environments.
+As confirmed by several industrial collaborators [7], Jupyter notebooks often serve as the starting point for developing ML applications before transitioning to production-ready ML pipelines, where much of the Python code used in notebooks is reused in large-scale distributed computing environments.
 This widespread adoption, even in mission-critical applications such as those at NASA Jet Propulsion Lab [8], underscores the importance of quality assurance for notebooks.
@@
 Finding 1.6: Name-related errors, in particular Variable Not Found, are mainly due to notebook-specific issues.
-To assess the accuracy of our root cause analysis, we try to reproduce 30 crashes in Kaggle and successfully reproduce 20.
-We then resolve the crashes, finding that 18 ($90\%$) are caused by the root causes identified in this study.
-This high accuracy suggests that our approach is reliable.
+Evaluating root cause accuracy
+To assess the accuracy of our root cause analysis, we randomly selected 30 crashes in Kaggle notebooks and attempted to reproduce and fix them.
+For each case, we download the input datasets (if available), execute a subset of code cells necessary to trigger the crash, localize and repair the faulty code to verify if the labeled root cause aligns with the actual fault.
+A crash is considered fixed if the previously crashing cell executes successfully while preserving the original intent.
+We successfully reproduced and fixed 20 out of the 30 crashes.
+Of the 10 unsuccessful cases, 8 lack input datasets, and the remaining two no longer crash.
+From the 20 cases, 18 ($90\%$) match the root cause identified by this study, which suggests that our approach is reliable.
+For the two misclassifications, one involves an index error due to out-of-order execution mislabeled as API misuse.
+The other is a tensor shape mismatch caused by faulty logic (i.e., implementation error) in earlier cells, incorrectly labeled as data confusion.
+The main challenge in labeling errors and root causes without executing the notebooks lies in mentally tracking the internal kernel state.
+This can lead to misunderstandings by labelers, especially when the code logic is complex.
 Figure 4: Overview of the co-occurrence of crash types and root causes (see Table V for full names).
@@
 Future work could address this by analyzing version histories or autosaved checkpoints to capture a more complete development process.
+Furthermore, although we have filtered out tutorial notebooks, the ML notebooks in our dataset may still not fully represent those used as prototypes in an industrial setting.
+This introduces a potential threat to validity when generalizing crash patterns that could propagate from notebooks to production.
+Future work should address this by studying ML prototype notebooks developed within industry to better assess how our findings apply.
 Temporal limitations in datasets
@@
 We manually label ML notebook crashes, which can be prone to errors.
-To enhance the classification reliability, we use a structured two-stage process.
+To enhance the reliability of classification, we use a structured two-stage process.
 In the labeling stage, three evaluators independently and iteratively categorize crashes, followed by discussions to synchronize findings, reduce bias, and promote consistency.
 In the reviewing stage, a reviewer validates the labels.
-Any disagreements are resolved through consensus.
+Any disagreements are resolved by consensus.
 Reproducibility of crashes and root cause analysis
-The error outputs we analyze represent real execution crashes, but they are not necessarily reproducible due to factors like execution order, missing datasets, or environment replication [2].
+The error outputs we analyze represent real execution crashes, but they are not necessarily reproducible due to factors such as execution order, missing datasets, or environment replication [2].
 While reproducing crashes could further validate our findings, it is challenging, especially for GitHub notebooks.